### v1

# Helmlab: A Data-Driven Analytical Color Space for UI Design Systems

Abstract
We present Helmlab, a 72-parameter analytical color space for UI design systems. The forward transform maps CIE XYZ to a perceptually-organized Lab representation through learned matrices, per-channel power compression, Fourier hue correction, and embedded Helmholtz–Kohlrausch lightness adjustment. A post-pipeline neutral correction guarantees that achromatic colors map to a=b=0a\!=\!b\!=\!0 (chroma $10^{-6}$), and a rigid rotation of the chromatic plane improves hue-angle alignment without affecting the distance metric, which is invariant under isometries. On the combvd dataset (3,813 color pairs), Helmlab achieves a stress of 23.22, a 20.4% reduction from CIEDE2000 (29.18). Cross-validation on He et al. 2022 and MacAdam 1974 shows competitive cross-dataset performance. The transform is invertible with round-trip errors below $10^{-14}$. Gamut mapping, design-token export, and dark/light mode adaptation utilities are included for use in web and mobile design systems.
Keywords: color space, color difference, perceptual uniformity, STRESS, Helmholtz–Kohlrausch, UI design systems
1 Introduction
UI design systems need accurate perceptual distance prediction for contrast checking, an achromatic axis where grays map to zero chroma for gradient interpolation, and reasonable hue-angle alignment for programmatic color manipulation. No existing color space provides all three.
CIE Lab [1] and CIEDE2000 [2] are the standard for color-difference evaluation, but their stress on combvd is 29.18. Oklab [3] is simple and part of CSS Color Level 4, but was optimized for hue uniformity rather than color-difference prediction, and gets stress 47.5 (Euclidean) on combvd. CAM16-UCS [4] is derived from a color appearance model but is computationally expensive and has achromatic chroma leakage ($\bar{C}=1.42$). IPT [5] was optimized for hue linearity at the expense of lightness accuracy. $J_{z}a_{z}b_{z}$ [6] targets HDR via the PQ curve but is complex and not CSS-native. None of these spaces jointly optimize the color-space transform and distance metric end-to-end against psychophysical data.
In this paper, we treat the entire XYZ-to-distance pipeline as a single system with 72 learnable parameters and optimize all of them jointly against the largest available color-difference dataset. The contributions are:
  1. 1.
    A 72-parameter analytical color space with stress 23.22 on combvd, 20.4% lower than CIEDE2000.
  2. 2.
    A neutral correction that guarantees $C<10^{-6}$ for achromatic colors, at a stress cost of only +0.04+0.04.
  3. 3.
    A rigid rotation in the $ab$ plane that reduces hue error (RMS 16.1°) without changing the distance metric (the metric is invariant under rotation).
  4. 4.
    Embedded Helmholtz–Kohlrausch correction where lightness depends on chroma, with parameters learned from data.
  5. 5.
    Gamut mapping, CSS/Android/iOS token export, and dark/light mode adaptation for design system use.
2 Design Goals and Metrics
Helmlab is optimized for UI design systems operating in sRGB and Display P3 gamuts under typical viewing conditions (D65 illuminant, average surround). We define five quantitative targets:
Color-difference accuracy (stress).
The STRESS metric [7] measures agreement between predicted distances $\Delta E_{i}$ and visual differences $DV_{i}$:
$\mathrm{STRESS}=100\sqrt{\frac{\sum_{i}(DV_{i}-F\cdot\Delta E_{i})^{2}}{\sum_{i}DV_{i}^{2}}}$ (1)
where $F=\sum DV_{i}\Delta E_{i}/\sum\Delta E_{i}^{2}$. Lower is better. We train on combvd (3,813 pairs from six experiments [7]) and cross-validate on He et al. 2022 [8] (82 pairs) and MacAdam 1974 [9] (128 pairs).
Achromatic integrity.
Neutral colors (D65-adapted grays, i.e.,  stimuli on the achromatic axis proportional to the D65 white point) must map to $a=b=0$. We measure maximum chroma $C_{\max}$ over 256 gray levels spanning $Y\in[0.001,2.0]$.
Hue-angle alignment.
[…]
Round-trip accuracy.
For any sRGB color, $\|T^{-1}(T(\mathbf{x}))-\mathbf{x}\|_{\infty}$ should be below $10^{-8}$.
Munsell uniformity.
Coefficient of variation (CV) of neighbor-pair distances along Munsell constant-hue, constant-value lines (3,590 pairs). Lower CV indicates more uniform spacing.
3 The Helmlab Color Space
3.1 Forward Transform
The forward transform maps CIE XYZ (D65) to Lab through eleven stages (Figure˜1). All parameters are jointly optimized (Section˜5).
Figure 1: The Helmlab forward transform pipeline. Blue: linear operations; yellow: nonlinear corrections; green: structural guarantees (NC, rotation). All 72 parameters are jointly optimized.
Stage 1: Cone response (9 params).
A learned $3\!\times\!3$ matrix $\mathbf{M}_{1}$ maps XYZ to cone-response-like signals:
$\mathbf{c}=\mathbf{M}_{1}\begin{bmatrix}X\\ Y\\ Z\end{bmatrix}$ (2)
Unlike fixed matrices (Hunt–Pointer–Estévez, Bradford), $\mathbf{M}_{1}$ is freely optimized.
Stage 2: Power compression (3 params).
Each channel undergoes signed power compression:
$c_{i}^{\prime}=\mathrm{sign}(c_{i})\,|c_{i}|^{\gamma_{i}},\quad i\in\{0,1,2\}$ (3)
The exponents $\gamma_{i}$ generalize the cube-root of CIE Lab ($\gamma=1/3$). The optimized values are close to 0.4, which is between the cube-root (0.33) and square-root (0.5) compressions used in other spaces.
Stage 3: Lab projection (9 params).
A second matrix $\mathbf{M}_{2}$ projects compressed signals to opponent channels:
$\begin{bmatrix}L_{\mathrm{raw}}\\ a_{\mathrm{raw}}\\ b_{\mathrm{raw}}\end{bmatrix}=\mathbf{M}_{2}\,\mathbf{c}^{\prime}$ (4)
Stage 4: Hue correction (8 params).
A 4-harmonic Fourier rotation corrects systematic hue distortions:
$\delta(h)=\sum_{k=1}^{4}\!\bigl[\alpha_{k}\cos(kh)+\beta_{k}\sin(kh)\bigr]$ (5)
where $h=\mathrm{atan2}(b,a)$. The chromatic vector $(a,b)$ is rotated by $\delta(h)$, preserving chroma while redistributing hue angles.
Stage 5: Helmholtz–Kohlrausch embedding (6 params).
Saturated colors appear brighter than achromatic stimuli of equal luminance [1]. We embed this effect in the lightness channel directly:
$L\mathrel{+}=w_{\mathrm{HK}}\cdot C^{p_{\mathrm{HK}}}\cdot\bigl[1+f_{\mathrm{HK}}(h)\bigr]$ (6)
where $C=\sqrt{a^{2}+b^{2}}$ and $f_{\mathrm{HK}}(h)$ is a 2-harmonic hue modulation. The optimized weight is $w_{\mathrm{HK}}=0.37$ and the power is $p_{\mathrm{HK}}=0.85$, making lightness explicitly dependent on chroma. Most existing spaces treat lightness and chroma as independent.
Stage 6: Lightness refinement (8 params).
A cubic polynomial (3 params) with hue-dependent additive term (2 params), followed by dark-region exponential compression (3 params):
$L_{1}$ $=p_{1}L^{3}+p_{2}L^{2}+p_{3}L+t\cdot f_{\mathrm{Lh}}(h)$ (7)
$L_{2}$ $=L_{1}\cdot\exp\!\bigl[\lambda_{d}(h)\cdot L_{1}(1-L_{1})^{2}\bigr]$ (8)
where t=L(1L)t=L(1-L) and $f_{\mathrm{Lh}}(h)$ is a 1-harmonic Fourier correction targeting the hue–lightness interaction sign flip (dark cyans vs mid cyans). The dark compression coefficient $\lambda_{d}$ includes hue modulation to handle the known difference in dark-blue vs dark-yellow perception.
Stages 7–8: Chroma processing (18 params).
Chroma processing proceeds in four interleaved sub-stages: (a) hue-dependent scaling (4 harmonics, 8 params):
$(a,b)\mathrel{*}=\exp\!\Bigl[\sum_{k=1}^{4}\!\bigl[\alpha^{\prime}_{k}\cos(kh)+\beta^{\prime}_{k}\sin(kh)\bigr]\Bigr]$ (9)
[…]
Stage 9: Hue-dependent lightness (4 params).
A final $L\mathrel{*}=\exp[g(h)]$ captures residual hue–lightness interactions (e.g.,  blue appears darker, yellow lighter at equal luminance).
Stage 10: Neutral correction.
With independent $\gamma_{i}$ per channel, the D65 achromatic axis (stimuli proportional to $X_{D65},Y_{D65},Z_{D65}]$) generally does not map to a=b=0a\!=\!b\!=\!0 after the enrichment stages. We correct this by running 256 gray levels ($Y\in[0.001,2.0]$) through stages 1–9, recording the achromatic error $a_{\mathrm{err}}(L)$, $b_{\mathrm{err}}(L)$, and fitting PCHIP interpolants. At runtime:
$a\leftarrow a-a_{\mathrm{err}}(L),\quad b\leftarrow b-b_{\mathrm{err}}(L)$ (10)
Because the correction is applied after all nonlinear stages, it accounts for every distortion in the pipeline. The resulting maximum chroma for grays is $10^{-6}$, limited only by PCHIP interpolation error on the 256-point LUT (the underlying achromatic curve is smooth, so interpolation error is negligible).
Stage 11: Rigid rotation ($\varphi$).
A uniform rotation of the $(a,b)$ plane by angle $\varphi$:
$\begin{bmatrix}a^{\prime}\\ b^{\prime}\end{bmatrix}=\begin{bmatrix}\cos\varphi&-\sin\varphi\\ \sin\varphi&\cos\varphi\end{bmatrix}\begin{bmatrix}a\\ b\end{bmatrix}$ (11)
This is an isometry: it preserves $\Delta a^{2}+\Delta b^{2}$ for any pair, and therefore preserves the distance metric exactly (Section˜3.2). The rotation angle $\varphi=-28.2$° is chosen to minimize the maximum hue-angle error across sRGB primaries and secondaries.
3.2 Rotation Invariance
The distance metric (Section˜4) depends on $\Delta L$, $\Delta a^{2}+\Delta b^{2}$, $\bar{L}$, and $\bar{C}=\sqrt{\bar{a}^{2}+\bar{b}^{2}}$. A rigid rotation of the $(a,b)$ plane preserves each of these quantities: $\Delta a^{\prime 2}+\Delta b^{\prime 2}=\Delta a^{2}+\Delta b^{2}$ and $\bar{a}^{\prime 2}+\bar{b}^{\prime 2}=\bar{a}^{2}+\bar{b}^{2}$. Therefore the distance is exactly invariant, and we verified this empirically: the COMBVD stress difference before and after rotation is $0.0000000000$.
This means any hue-angle improvement from the rotation is obtained at zero cost to color-difference prediction.
3.3 Surround Parameter
Helmlab includes provisions for a surround luminance parameter $S\in[0,1]$ (0=dark, 0.5=average, 1=bright) that modulates several pipeline stages: H-K weight, dark compression, chroma scaling, and L-dependent chroma. In the current release, the surround parameters are set to zero (trained on average-surround data only). The architecture is ready for future training on viewing-condition-dependent datasets. The UI layer uses $S$ for heuristic dark/light mode adaptation via soft L-inversion.
3.4 Inverse Transform
Each stage has an explicit inverse. Matrix stages invert via $\mathbf{M}_{1}^{-1}$, $\mathbf{M}_{2}^{-1}$; power compression via $c_{i}=\mathrm{sign}(c_{i}^{\prime})|c_{i}^{\prime}|^{1/\gamma_{i}}$ (closed form). Hue corrections and lightness corrections are inverted via Newton iteration (3–5 iterations, converging to machine precision). The neutral correction inverts by adding back the PCHIP-interpolated error. Round-trip accuracy across the full sRGB gamut is $10^{-14}$.
4 Distance Metric
Given two colors with coordinates $(L_{1},a_{1},b_{1})$ and $(L_{2},a_{2},b_{2})$, we compute:
$S_{L}$ $=1+s_{L}(\bar{L}-0.5)^{2}$ (12)
$S_{C}$ $=1+s_{C}\bar{C}$ (13)
where $\bar{L}$, $\bar{C}$ are pair averages. The raw distance is:
$d=\left[\left(\frac{\Delta L}{S_{L}}\right)^{\!2}+w_{C}\frac{\Delta a^{2}+\Delta b^{2}}{S_{C}^{2}}\right]^{p/2}$ (14)
followed by monotonic compression and post-power:
$\Delta E=\left[\frac{d}{1+c\cdot d}\right]^{q}$ (15)
The pair-dependent weights $S_{L}$ and $S_{C}$ are inspired by CIEDE2000 but learned from data (optimized values: $s_{L}\approx 10^{-4}$, $s_{C}\approx 0.0045$). The Minkowski exponent $p=0.83$ and chroma weight $w_{C}=1.03$ slightly favor chroma differences over lightness. The compression $c=1.59$ and post-power $q=1.1$ reduce systematic bias for large color differences.
The distance metric contributes 7 parameters. Total system: 9+3+9+8+6+8+18+4 = 65 space + 7 metric = 72 parameters.
5 Optimization
All 72 parameters are jointly optimized by minimizing:
$\mathcal{L}=\mathrm{STRESS}_{\textsc{combvd}}+0.05\cdot\mathrm{STRESS}_{\mathrm{He}}$ (16)
using L-BFGS-B [10] with box constraints. We perform 8 random restarts (each 5,000 iterations) and keep the best. The He et al. term ($\lambda=0.05$) acts as a mild regularizer against overfitting to combvd.
Training data.
combvd comprises 3,813 color pairs from six psychophysical experiments: BFD-P(D65) (2,028), BFD-P(M) (548), BFD-P(C) (200), Witt (418), RIT-DuPont (312), Leeds (307). All pairs have visual-difference values on a common scale. We train on the full dataset (no train/val split for the final model), monitoring a held-out 20% split (seed=42) during development to check for overfitting.
Cross-validation data.
He et al. 2022 (82 pairs, 10° observer) enters the loss as a regularizer ($\lambda\!=\!0.05$) and is therefore not fully independent; its stress is reported for transparency but should be interpreted as a semi-held-out metric. MacAdam 1974 (128 pairs, 10°) is held out entirely and never seen during optimization.
Convergence.
The optimizer typically converges within 3,000–5,000 function evaluations per restart. The validation gap (train–val stress difference) is +1.11, which shows some overfitting but not severe for a 72-parameter model.
6 Evaluation
6.1 Color-Difference Prediction
Figure˜2 and Table˜1 show stress on combvd for Helmlab and nine baseline methods. Helmlab achieves 23.22, compared to 29.18 for CIEDE2000 (-20.4%). A 10,000-iteration paired bootstrap gives Helmlab 95% CI $[22.50,23.93]$ and CIEDE2000 95% CI $[27.64,30.84]$. The intervals do not overlap ($p<10^{-4}$).
Figure 2: stress on COMBVD (3,813 pairs). Lower is better.
Table 1: stress on combvd (3,813 pairs). Lower is better.
Method stress vs CIEDE2000
Helmlab 23.22 -20.4%
CIEDE2000 29.18
$\Delta E^{*}_{ab}$ (CIE76) 42.80 +46.7%
CIE94 33.59 +15.1%
$\Delta E$ CMC 34.04 +16.6%
CAM16-UCS (Euclid.) 33.90 +16.2%
IPT (Euclid.) 41.21 +41.3%
$J_{z}a_{z}b_{z}$ (Euclid.) 41.90 +43.6%
Oklab (Euclid.) 47.46 +62.7%
sRGB (Euclid.) 67.82 +132.5%
Figure 3: Predicted vs observed color differences on COMBVD.
The baselines in Table˜1 use Euclidean distance in each space’s native Cartesian coordinates, without parametric weighting or compression. Oklab’s high stress (47.46) is expected because it was optimized for hue uniformity [3], not color-difference prediction on combvd. The Helmlab distance metric includes pair-dependent weighting ($S_{L}$, $S_{C}$), a Minkowski exponent, and monotonic compression, so the comparison is not on equal terms. With Euclidean distance only, Helmlab gets stress $\approx$30.2, which is close to CIEDE2000 (29.18). This means the space transform accounts for roughly half the improvement and the distance metric for the other half.
6.2 Cross-Validation
On He et al. 2022 (regularizer, 82 pairs), Helmlab achieves stress 29.0 vs CIEDE2000’s 32.6. On MacAdam 1974 (held out entirely, 128 pairs), Helmlab achieves 20.2 vs CAM16-UCS’s 18.7 (best). Figure˜4 visualizes these results. Helmlab outperforms CIEDE2000 on He et al. but trails CAM16-UCS on MacAdam. This is consistent with what one would expect from a model trained primarily on combvd: it generalizes reasonably but does not dominate on all evaluation sets.
Figure 4: Cross-dataset validation on held-out data.
6.3 Generation Properties: The Measurement–Generation Tradeoff
A color space optimized for distance prediction does not necessarily produce usable coordinates for color generation. During development, the measurement-optimal configuration mapped neutral grays to chroma $\bar{C}\approx 0.34$. The optimizer had deformed the achromatic axis to reduce stress, which made gray gradients unusable because they showed visible color artifacts. Adding a penalty to constrain grays to $C=0$ did not help: the optimizer either stayed in the measurement-optimal basin or escaped to a different basin with much worse stress (26–31). There was no smooth transition between the two.
The neutral correction (Section˜3, Stage 10) addresses this problem. It subtracts the achromatic error after all enrichment stages, so the optimizer can find its measurement-optimal parameters without constraint. The achromatic axis is then corrected separately. The stress increase from this correction is +0.04+0.04 (23.18 to 23.22), while achromatic chroma drops from 0.34 to $10^{-6}$.
Figure˜5 shows neutral ramp uniformity and achromatic chroma leakage. Helmlab’s neutral correction eliminates chroma leakage entirely ($C<10^{-6}$), while CAM16-UCS exhibits significant leakage ($\bar{C}=1.42$). Table˜2 summarizes all generation properties.
Figure 5: Left: neutral ramp step uniformity. Right: achromatic chroma leakage. Helmlab’s NC guarantees $C<10^{-6}$ for all grays.
Table 2: Generation quality metrics.
Property Value
Achromatic max $C$ $10^{-6}$
Hue RMS (6 primaries) 16.1°
Hue max error 20.2°
Round-trip $\|\cdot\|_{\infty}$ $10^{-14}$
Munsell CV 17.6%
Jacobian det (min) 0.10
Jacobian cond (median) 7.7
Val gap (train-val) +1.11
He stress 29.03
MacAdam stress 20.23
The achromatic guarantee means gray gradients and lightness ramps interpolate without color artifacts, which is necessary for design-token generation and palette export. The hue RMS of 16.1° means sRGB primaries land within about 20° of their expected positions. This is acceptable for palette generation but not perfect. The rigid rotation (Section˜3, Stage 11) provides this hue improvement without affecting the distance metric because the metric is invariant under rotation. The Jacobian was evaluated on a $64\!\times\!64\!\times\!64$ sRGB grid via finite differences; the minimum determinant (0.10) is positive throughout, so the map is locally invertible and orientation-preserving across the entire sRGB gamut.
6.4 Ablation Study
Table˜3 isolates the contribution of key components.
Table 3: Ablation: effect of removing components.
Configuration stress $\Delta$
Full Helmlab 23.22
Euclidean distance only $\sim$30.2 +7.0
No H-K embedding ($w_{\mathrm{HK}}=0$) $\sim$24.5 +1.3
No hue correction $\sim$25.8 +2.6
No dark L compression $\sim$23.9 +0.7
Rotation (-28.2°) 23.22 0.00
No rotation 23.22 0.00
With Euclidean distance, the space alone gets stress 30.2, which is slightly worse than CIEDE2000 (29.18). The full metric brings it down to 23.22, recovering the remaining 7 points. Compared to other Euclidean baselines (CAM16-UCS = 33.9), the space alone already gives a clear improvement. Among individual stages, the H-K embedding and hue correction contribute the most. The rotation row confirms that the metric is exactly invariant under rotation.
7 Practical Adoption
Figure˜6 shows the sRGB gamut boundary in Helmlab coordinates at three lightness levels. The boundaries are smooth and well-behaved. Helmlab includes a Python utility layer for design system integration:
Figure 6: sRGB gamut in Helmlab aabb plane at three L levels. Points colored by their sRGB values.
Gamut mapping.
Binary-search chroma reduction along the L axis maps out-of-gamut Helmlab coordinates to the nearest in-gamut sRGB or Display P3 color, preserving hue and lightness.
Contrast utilities.
ensure_contrast(fg, bg, min_ratio) adjusts foreground lightness via binary search on the Helmlab L axis to meet WCAG 2.x contrast requirements (4.5:1 for AA, 7:1 for AAA). Hue and chroma are preserved.
Palette generation.
Lightness ramps and hue rings at fixed chroma produce perceptually-spaced palettes. Semantic scales (Tailwind-style 50–950) are generated by interpolating L between 0.97 and 0.10 with chroma/hue preserved.
Dark/light mode.
When surround parameters are trained, the surround $S$ directly modulates the transform. Currently, a fallback soft L-inversion maps colors between light ($S=0.7$) and dark ($S=0.2$) modes with contrast enforcement.
Token export.
A single TokenExporter produces CSS custom properties (oklch(), color(display-p3 ...)), Android XML (using HCT tone/chroma/hue semantics for Material Design compatibility), iOS Swift (Display P3), Tailwind config, and raw JSON.
8 Limitations and Future Work
There are several limitations worth noting:
  1. 1.
    Training data. combvd uses a 2° standard observer, which may not generalize to large-field (10°) viewing. The MacAdam cross-validation partially addresses this.
  2. 2.
    Model complexity. 72 parameters is considerably more than Oklab ($\sim$15 designed parameters). Each stage has a perceptual motivation, but the overall model is harder to interpret than simpler spaces.
  3. 3.
    Hue alignment. RMS hue error of 16.1° is a compromise. Some hue regions (cyan, yellow) have errors up to 20°. Further harmonic terms could improve this at the risk of overfitting.
  4. 4.
    Surround dependency. The surround parameter $S$ is architecturally present but not yet trained on viewing-condition data. Current dark/light adaptation uses heuristics.
  5. 5.
    Sub-dataset variation. Helmlab underperforms CIEDE2000 on 3 of 6 combvd sub-datasets (BFD-P(C), Leeds, RIT-DuPont), which may indicate that these subsets benefit more from CIE Lab-based corrections.
  6. 6.
    Not general-purpose. Helmlab is optimized for UI workflows (sRGB/P3 gamut, screen viewing). It is not designed for print, photography, or spectral applications.
  7. 7.
    Statistical reporting. Bootstrap confidence intervals confirm the aggregate improvement over CIEDE2000 ($p<10^{-4}$), but per-sub-dataset significance and parameter uncertainty across restarts remain unquantified.
For future work, we plan to train the surround parameter on paired light/dark viewing data and to expand the training set with targeted human feedback using the bidirectional optimization framework that is already in the codebase.
9 Conclusion
We have presented Helmlab, a 72-parameter analytical color space that achieves stress 23.22 on combvd, which is 20.4% lower than CIEDE2000. The achromatic axis is guaranteed to have $C<10^{-6}$ via the neutral correction, and a rigid rotation brings hue error down to 16.1° RMS without affecting the distance metric. The neutral correction is the key element that makes the space usable for both measurement and generation: it lets the optimizer find its best parameters freely and then fixes the achromatic axis at a cost of only +0.04+0.04 stress points. Code, trained parameters, and utilities are available at https://github.com/Grkmyldz148/helmlab. An interactive demo is at https://grkmyldz148.github.io/helmlab/demo.html and documentation at https://grkmyldz148.github.io/helmlab/.
Acknowledgments
Thanks to the 71 anonymous observers who participated in the bidirectional evaluation study during development.
Appendix A Parameter Table
Table˜4 lists all 72 optimized parameters grouped by pipeline stage.
Table 4: Optimized Helmlab parameters (v19-NC, $\varphi=-28.2$°).
Stage Parameter Value Description
$\mathbf{M}_{1}$ (9) Row 0 [0.724, 0.238, -0.164] XYZ $\to$ cone-like
Row 1 [-0.325, 1.238, -0.024]
Row 2 [0.101, 0.341, 0.849]
$\gamma$ (3) $\gamma_{0},\gamma_{1},\gamma_{2}$ 0.395, 0.421, 0.434 Power compression
$\mathbf{M}_{2}$ (9) Row 0 [-0.425, 0.489, 0.684] Compressed $\to$ Lab
Row 1 [1.931, -2.728, 0.765]
Row 2 [0.009, 1.642, -1.250]
Hue corr. (8) $\alpha_{1..4},\beta_{1..4}$ see JSON 4-harmonic Fourier
H-K (6) $w,p,m,s_{1},c_{2},s_{2}$ 0.370, 0.852, -0.487, … Chroma $\to$ L
Light. (8) $p_{1},p_{2},p_{3},Lh_{c1},Lh_{s1},\lambda_{d},h_{c},h_{s}$ 0.225, -0.748, 0.699, … Cubic + Lh + dark
Chroma (18) 4-harm + power + $lc_{1},lc_{2}$ + HLC see JSON Hue/power/L-dep
Hue-L (4) $g_{c1..2},g_{s1..2}$ see JSON Hue $\to$ L
Distance (7) $s_{L}$ $9.16\times 10^{-5}$ Pair-dep L weight
$s_{C}$ 0.0045 Pair-dep C weight
pp 0.829 Minkowski exponent
$w_{C}$ 1.028 Chroma weight
$c$ 1.588 Compression
$q$ 1.100 Post-power
$\alpha$ 0.000 Linear term
Full parameter values are available in the accompanying analytical_params.json file.

### v2

@@
-Helmlab: A Data-Driven Analytical Color Space for UI Design Systems
+arXiv:2602.23010v2[cs.GR] 01 Mar 2026
 Abstract
@@
 A post-pipeline neutral correction guarantees that achromatic colors map to a=b=0a\!=\!b\!=\!0 (chroma $10^{-6}$), and a rigid rotation of the chromatic plane improves hue-angle alignment without affecting the distance metric, which is invariant under isometries.
-On the combvd dataset (3,813 color pairs), Helmlab achieves a stress of 23.22, a 20.4% reduction from CIEDE2000 (29.18).
+On the combvd dataset (3,813 color pairs), Helmlab achieves a stress of 23.30, a 20.2% reduction from CIEDE2000 (29.18).
+A blue-band refit with sub-dataset penalties reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ at a cost of only +0.08+0.08 stress.
 Cross-validation on He et al.
@@
   1. 1.
-    A 72-parameter analytical color space with stress 23.22 on combvd, 20.4% lower than CIEDE2000.
+    A 72-parameter analytical color space with stress 23.30 on combvd, 20.2% lower than CIEDE2000.
   2. 2.
@@
   3. 3.
-    A rigid rotation in the $ab$ plane that reduces hue error (RMS 16.1°) without changing the distance metric (the metric is invariant under rotation).
+    A rigid rotation in the $ab$ plane that reduces hue error (RMS 18.1°) without changing the distance metric (the metric is invariant under rotation).
   4. 4.
@@
   5. 5.
+    A blue-band refit that reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ while improving stress on 5 of 6 combvd sub-datasets.
+  6. 6.
     Gamut mapping, CSS/Android/iOS token export, and dark/light mode adaptation for design system use.
@@
 where $C=\sqrt{a^{2}+b^{2}}$ and $f_{\mathrm{HK}}(h)$ is a 2-harmonic hue modulation.
-The optimized weight is $w_{\mathrm{HK}}=0.37$ and the power is $p_{\mathrm{HK}}=0.85$, making lightness explicitly dependent on chroma.
+The optimized weight is $w_{\mathrm{HK}}=0.389$ and the power is $p_{\mathrm{HK}}=0.849$, making lightness explicitly dependent on chroma.
 Most existing spaces treat lightness and chroma as independent.
@@
 $\Delta E=\left[\frac{d}{1+c\cdot d}\right]^{q}$ (15)
-The pair-dependent weights $S_{L}$ and $S_{C}$ are inspired by CIEDE2000 but learned from data (optimized values: $s_{L}\approx 10^{-4}$, $s_{C}\approx 0.0045$).
-The Minkowski exponent $p=0.83$ and chroma weight $w_{C}=1.03$ slightly favor chroma differences over lightness.
-The compression $c=1.59$ and post-power $q=1.1$ reduce systematic bias for large color differences.
+The pair-dependent weights $S_{L}$ and $S_{C}$ are inspired by CIEDE2000 but learned from data (optimized values: $s_{L}=0.001$, $s_{C}=0.022$).
+The Minkowski exponent $p=0.804$ and chroma weight $w_{C}=1.046$ slightly favor chroma differences over lightness.
+The compression $c=1.590$ and post-power $q=1.1$ reduce systematic bias for large color differences.
 The distance metric contributes 7 parameters. Total system: 9+3+9+8+6+8+18+4 = 65 space + 7 metric = 72 parameters.
@@
 All 72 parameters are jointly optimized by minimizing:
-$\mathcal{L}=\mathrm{STRESS}_{\textsc{combvd}}+0.05\cdot\mathrm{STRESS}_{\mathrm{He}}$ (16)
+$\mathcal{L}=\mathrm{STRESS}_{\textsc{combvd}}+0.05\cdot\mathrm{STRESS}_{\mathrm{He}}+\mathcal{L}_{\mathrm{blue}}$ (16)
 using L-BFGS-B [10] with box constraints.
-We perform 8 random restarts (each 5,000 iterations) and keep the best.
+We perform 10 random restarts (each 5,000 iterations) and keep the best.
 The He et al.
 term ($\lambda=0.05$) acts as a mild regularizer against overfitting to combvd.
+The blue-band penalty is:
+$\mathcal{L}_{\mathrm{blue}}=\lambda_{b}(\mathrm{STRESS}_{\mathrm{RIT}}+\mathrm{STRESS}_{\mathrm{LEEDS}}+0.5\,\mathrm{STRESS}_{\mathrm{FullBlue}})$ (17)
+where $\lambda_{b}=0.1$ and the subscripts denote combvd sub-datasets with high blue-band content.
+Additional constraints enforce round-trip accuracy ($\lambda_{\mathrm{RT}}=100$), achromatic integrity ($\lambda_{\mathrm{ach}}=500$), and Munsell uniformity ($0.1\cdot\max(\mathrm{CV}-20,0)$).
+The blue-band penalty was introduced in v2 of this paper to address gradient non-uniformity in the blue–cyan region (Section˜6.6).
 Training data.
@@
 The optimizer typically converges within 3,000–5,000 function evaluations per restart.
-The validation gap (train–val stress difference) is +1.11, which shows some overfitting but not severe for a 72-parameter model.
+Five-fold cross-validation (with re-optimization per fold) yields a mean validation gap of +0.98+0.98 stress points (std.
+$0.78$), indicating mild but acceptable overfitting for a 72-parameter model.
 6 Evaluation
@@
 Figure˜2 and Table˜1 show stress on combvd for Helmlab and nine baseline methods.
-Helmlab achieves 23.22, compared to 29.18 for CIEDE2000 (-20.4%).
+Helmlab achieves 23.30, compared to 29.18 for CIEDE2000 (-20.2%).
 A 10,000-iteration paired bootstrap gives Helmlab 95% CI $[22.50,23.93]$ and CIEDE2000 95% CI $[27.64,30.84]$.
@@
 Method stress vs CIEDE2000
-Helmlab 23.22 -20.4%
+Helmlab 23.30 -20.2%
 CIEDE2000 29.18
@@
 The Helmlab distance metric includes pair-dependent weighting ($S_{L}$, $S_{C}$), a Minkowski exponent, and monotonic compression, so the comparison is not on equal terms.
-With Euclidean distance only, Helmlab gets stress $\approx$30.2, which is close to CIEDE2000 (29.18).
-This means the space transform accounts for roughly half the improvement and the distance metric for the other half.
+With Euclidean distance only, Helmlab gets stress 27.6, which already improves on CIEDE2000 (29.18) by 5.4%.
+The full distance metric then reduces stress by a further 4.3 points.
+Relative to CIEDE2000, the space transform and the distance metric contribute roughly equally to the total improvement.
 6.2 Cross-Validation
 On He et al.
-2022 (regularizer, 82 pairs), Helmlab achieves stress 29.0 vs CIEDE2000’s 32.6.
-On MacAdam 1974 (held out entirely, 128 pairs), Helmlab achieves 20.2 vs CAM16-UCS’s 18.7 (best).
+2022 (regularizer, 82 pairs), Helmlab achieves stress 30.3 vs CIEDE2000’s 32.6.
+On MacAdam 1974 (held out entirely, 128 pairs), Helmlab achieves 20.3 vs CAM16-UCS’s 18.7 (best).
 Figure˜4 visualizes these results.
@@
 The achromatic axis is then corrected separately.
-The stress increase from this correction is +0.04+0.04 (23.18 to 23.22), while achromatic chroma drops from 0.34 to $10^{-6}$.
+The stress increase from the neutral correction alone is +0.04+0.04 (from a measurement-only baseline of 23.18 to 23.22); the subsequent blue-band refit (Section˜6.6) adds another +0.08+0.08, yielding the final 23.30.
+Achromatic chroma drops from 0.34 to $10^{-6}$.
 Figure˜5 shows neutral ramp uniformity and achromatic chroma leakage.
@@
 Achromatic max $C$ $10^{-6}$
-Hue RMS (6 primaries) 16.1°
-Hue max error 20.2°
+Hue RMS (6 primaries) 18.1°
+Hue max error 23.9°
 Round-trip $\|\cdot\|_{\infty}$ $10^{-14}$
-Munsell CV 17.6%
+Munsell CV 20.0%
 Jacobian det (min) 0.10
 Jacobian cond (median) 7.7
-Val gap (train-val) +1.11
-He stress 29.03
-MacAdam stress 20.23
+5-fold CV gap (mean) +0.98
+He stress 30.33
+MacAdam stress 20.33
 The achromatic guarantee means gray gradients and lightness ramps interpolate without color artifacts, which is necessary for design-token generation and palette export.
-The hue RMS of 16.1° means sRGB primaries land within about 20° of their expected positions.
+The hue RMS of 18.1° means sRGB primaries land within about 24° of their expected positions.
 This is acceptable for palette generation but not perfect.
@@
 6.4 Ablation Study
-Table˜3 isolates the contribution of key components.
-Table 3: Ablation: effect of removing components.
+Table˜3 isolates the contribution of key components using frozen ablation: each row disables one component while keeping all other parameters at their optimized values.
+This measures the marginal contribution of each component given the current parameter setting, not its isolated effect.
+Table 3: Ablation: effect of removing components (frozen, not re-optimized).
 Configuration stress $\Delta$
-Full Helmlab 23.22
-Euclidean distance only $\sim$30.2 +7.0
-No H-K embedding ($w_{\mathrm{HK}}=0$) $\sim$24.5 +1.3
-No hue correction $\sim$25.8 +2.6
-No dark L compression $\sim$23.9 +0.7
-Rotation (-28.2°) 23.22 0.00
-No rotation 23.22 0.00
-With Euclidean distance, the space alone gets stress 30.2, which is slightly worse than CIEDE2000 (29.18).
-The full metric brings it down to 23.22, recovering the remaining 7 points.
-Compared to other Euclidean baselines (CAM16-UCS = 33.9), the space alone already gives a clear improvement.
-Among individual stages, the H-K embedding and hue correction contribute the most.
+Full Helmlab 23.30
+Euclidean distance only 27.60 +4.3
+No H-K embedding ($w_{\mathrm{HK}}=0$) 27.30 +4.0
+No hue correction 38.81 +15.5
+No dark L compression 23.49 +0.2
+Rotation (-28.2°) 23.30 0.00
+No rotation 23.30 0.00
+With Euclidean distance, the space alone gets stress 27.6, which already improves on CIEDE2000 (29.18) by 5.4%.
+The full metric brings it down to 23.30, recovering a further 4.3 points.
+Compared to other Euclidean baselines (CAM16-UCS = 33.9), the space alone gives a substantial improvement.
+Among individual stages, hue correction is the most critical ($\Delta=+15.5$), followed by the H-K embedding ($\Delta=+4.0$).
+These large frozen-ablation deltas reflect the co-optimization of all parameters: disabling one stage disrupts the balance of the entire pipeline.
 The rotation row confirms that the metric is exactly invariant under rotation.
+6.5 Sub-Dataset Performance
+Table˜4 breaks down stress by combvd sub-dataset.
+Helmlab outperforms CIEDE2000 on 3 of 6 sub-datasets (BFD-P(D65), BFD-P(M), Witt).
+On the blue-heavy sub-datasets (LEEDS, RIT-DuPont), the blue-band refit reduces the gap substantially compared to an earlier version without the penalty: RIT-DuPont improves from 23.49 to 20.41, nearly matching CIEDE2000’s 19.51.
+BFD-P(C) is the only sub-dataset measured under illuminant C (rather than D65) and with a 2° observer, which may explain the remaining deficit.
+Table 4: stress on individual combvd sub-datasets.
+Sub-dataset NN Helmlab CIEDE2000
+BFD-P(D65) 2,028 23.11 24.14
+BFD-P(M) 548 22.22 35.09
+BFD-P(C) 200 30.15 30.05
+Witt 418 28.88 30.37
+Leeds 307 20.92 19.46
+RIT-DuPont 312 20.41 19.51
+6.6 Blue-Band Analysis
+During development, we observed that an initial version of Helmlab (without the blue-band penalty) exhibited severe gradient non-uniformity in the blue–cyan region.
+A red-to-blue gradient interpolated in Helmlab Lab coordinates showed a $51.4\times$ max/min step-size ratio when measured with CIEDE2000, compared to $1.3\times$ for the same gradient in CIE Lab.
+Frozen ablation traced the problem to the 4-harmonic hue correction, whose Fourier coefficients overshoot in the blue–magenta band ($h\approx 240$340340°).
+Disabling hue correction reduced the gradient ratio from $51.4\times$ to $4.9\times$ but increased combvd stress by +15.5+15.5 points, making this an unacceptable trade-off.
+The solution was to add explicit sub-dataset penalties to the loss function (Equation˜17) and re-optimize all 72 parameters.
+With $\lambda_{b}=0.1$, the red-to-blue gradient ratio drops from $51.4\times$ to $5.8\times$ ($8.9\times$ improvement), while combvd stress increases by only +0.08+0.08 (from 23.22 to 23.30).
+stress improves on 5 of 6 sub-datasets (Table˜4), with the largest gain on RIT-DuPont (3.08-3.08 points).
+The only regression is BFD-P(D65) (+0.49+0.49), the largest sub-dataset, which accounts for the slight overall stress increase.
 7 Practical Adoption
@@
   1. 1.
-    Training data. combvd uses a 2° standard observer, which may not generalize to large-field (10°) viewing. The MacAdam cross-validation partially addresses this.
+    Training data.
+combvd predominantly uses the CIE 1964 10° standard observer ($\sim$95% of pairs).
+Only BFD-P(C) (200 pairs, 5%) uses the 2° observer.
+Gao et al.
+2023 [11] showed that the difference between 2° and 10° observers is not statistically significant for stress evaluation, and MacAdam 1974 (measured with the 2° observer) cross-validation confirms reasonable generalization.
   2. 2.
@@
   3. 3.
-    Hue alignment. RMS hue error of 16.1° is a compromise. Some hue regions (cyan, yellow) have errors up to 20°. Further harmonic terms could improve this at the risk of overfitting.
+    Hue alignment. RMS hue error of 18.1° is a compromise. Some hue regions (cyan) have errors up to 23.9°. Further harmonic terms could improve this at the risk of overfitting.
   4. 4.
@@
     Sub-dataset variation.
-Helmlab underperforms CIEDE2000 on 3 of 6 combvd sub-datasets (BFD-P(C), Leeds, RIT-DuPont), which may indicate that these subsets benefit more from CIE Lab-based corrections.
+Helmlab underperforms CIEDE2000 on 3 of 6 combvd sub-datasets (Table˜4).
+The blue-band refit improves 5 of 6 sub-datasets compared to the previous version; only BFD-P(C), measured under illuminant C with a 2° observer, remains notably worse (30.15 vs 30.05).
   6. 6.
@@
   7. 7.
+    Blue-band trade-off.
+The blue-band refit (Section˜6.6) reduces gradient non-uniformity by $8.9\times$ at a cost of +0.08+0.08 stress on combvd.
+The Munsell CV increases from 17.6% to 20.0%, hue RMS from 16.1° to 18.1°, and the He et al.
+cross-validation stress increases from 29.0 to 30.3.
+  8. 8.
     Statistical reporting.
 Bootstrap confidence intervals confirm the aggregate improvement over CIEDE2000 ($p<10^{-4}$), but per-sub-dataset significance and parameter uncertainty across restarts remain unquantified.
+The open-source release includes a second, generation-optimized space (GenSpace) with shared $\gamma\!=\!\tfrac{1}{3}$ and no enrichment stages, which trades distance accuracy for faster, invertible gradient and palette generation.
 For future work, we plan to train the surround parameter on paired light/dark viewing data and to expand the training set with targeted human feedback using the bidirectional optimization framework that is already in the codebase.
 9 Conclusion
-We have presented Helmlab, a 72-parameter analytical color space that achieves stress 23.22 on combvd, which is 20.4% lower than CIEDE2000.
-The achromatic axis is guaranteed to have $C<10^{-6}$ via the neutral correction, and a rigid rotation brings hue error down to 16.1° RMS without affecting the distance metric.
-The neutral correction is the key element that makes the space usable for both measurement and generation: it lets the optimizer find its best parameters freely and then fixes the achromatic axis at a cost of only +0.04+0.04 stress points.
+We have presented Helmlab, a 72-parameter analytical color space that achieves stress 23.30 on combvd, which is 20.2% lower than CIEDE2000.
+The achromatic axis is guaranteed to have $C<10^{-6}$ via the neutral correction, and a rigid rotation brings hue error down to 18.1° RMS without affecting the distance metric.
+A blue-band refit with sub-dataset penalties reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ at a cost of only +0.08+0.08 stress points, while improving stress on 5 of 6 sub-datasets.
+The neutral correction is the key element that makes the space usable for both measurement and generation: it lets the optimizer find its best parameters freely and then fixes the achromatic axis.
 Code, trained parameters, and utilities are available at https://github.com/Grkmyldz148/helmlab.
@@
 Appendix A Parameter Table
-Table˜4 lists all 72 optimized parameters grouped by pipeline stage.
-Table 4: Optimized Helmlab parameters (v19-NC, $\varphi=-28.2$°).
+Table˜5 lists all 72 optimized parameters grouped by pipeline stage.
+Table 5: Optimized Helmlab parameters (v20b, $\varphi=-28.2$°).
 Stage Parameter Value Description
-$\mathbf{M}_{1}$ (9) Row 0 [0.724, 0.238, -0.164] XYZ $\to$ cone-like
-Row 1 [-0.325, 1.238, -0.024]
-Row 2 [0.101, 0.341, 0.849]
-$\gamma$ (3) $\gamma_{0},\gamma_{1},\gamma_{2}$ 0.395, 0.421, 0.434 Power compression
-$\mathbf{M}_{2}$ (9) Row 0 [-0.425, 0.489, 0.684] Compressed $\to$ Lab
-Row 1 [1.931, -2.728, 0.765]
-Row 2 [0.009, 1.642, -1.250]
+$\mathbf{M}_{1}$ (9) Row 0 [0.734, 0.240, -0.158] XYZ $\to$ cone-like
+Row 1 [-0.330, 1.235, -0.000]
+Row 2 [0.086, 0.341, 0.832]
+$\gamma$ (3) $\gamma_{0},\gamma_{1},\gamma_{2}$ 0.389, 0.416, 0.424 Power compression
+$\mathbf{M}_{2}$ (9) Row 0 [-0.420, 0.485, 0.642] Compressed $\to$ Lab
+Row 1 [1.933, -2.718, 0.764]
+Row 2 [0.006, 1.628, -1.261]
 Hue corr. (8) $\alpha_{1..4},\beta_{1..4}$ see JSON 4-harmonic Fourier
-H-K (6) $w,p,m,s_{1},c_{2},s_{2}$ 0.370, 0.852, -0.487, … Chroma $\to$ L
-Light. (8) $p_{1},p_{2},p_{3},Lh_{c1},Lh_{s1},\lambda_{d},h_{c},h_{s}$ 0.225, -0.748, 0.699, … Cubic + Lh + dark
+H-K (6) $w,p,m,s_{1},c_{2},s_{2}$ 0.389, 0.849, -0.494, … Chroma $\to$ L
+Light. (8) $p_{1},p_{2},p_{3},Lh_{c1},Lh_{s1},\lambda_{d},h_{c},h_{s}$ 0.222, -0.727, 0.695, … Cubic + Lh + dark
 Chroma (18) 4-harm + power + $lc_{1},lc_{2}$ + HLC see JSON Hue/power/L-dep
 Hue-L (4) $g_{c1..2},g_{s1..2}$ see JSON Hue $\to$ L
-Distance (7) $s_{L}$ $9.16\times 10^{-5}$ Pair-dep L weight
-$s_{C}$ 0.0045 Pair-dep C weight
-pp 0.829 Minkowski exponent
-$w_{C}$ 1.028 Chroma weight
-$c$ 1.588 Compression
+Distance (7) $s_{L}$ $1.01\times 10^{-3}$ Pair-dep L weight
+$s_{C}$ 0.022 Pair-dep C weight
+pp 0.804 Minkowski exponent
+$w_{C}$ 1.046 Chroma weight
+$c$ 1.590 Compression
 $q$ 1.100 Post-power
 $\alpha$ 0.000 Linear term
-Full parameter values are available in the accompanying analytical_params.json file.
+Full parameter values are available in the accompanying metric_params.json file.

### v3

@@
-arXiv:2602.23010v2[cs.GR] 01 Mar 2026
+Helmlab: A Data-Driven Analytical Color Space for UI Design Systems
 Abstract
-We present Helmlab, a 72-parameter analytical color space for UI design systems.
-The forward transform maps CIE XYZ to a perceptually-organized Lab representation through learned matrices, per-channel power compression, Fourier hue correction, and embedded Helmholtz–Kohlrausch lightness adjustment.
-A post-pipeline neutral correction guarantees that achromatic colors map to a=b=0a\!=\!b\!=\!0 (chroma $10^{-6}$), and a rigid rotation of the chromatic plane improves hue-angle alignment without affecting the distance metric, which is invariant under isometries.
-On the combvd dataset (3,813 color pairs), Helmlab achieves a stress of 23.30, a 20.2% reduction from CIEDE2000 (29.18).
-A blue-band refit with sub-dataset penalties reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ at a cost of only +0.08+0.08 stress.
-Cross-validation on He et al.
-2022 and MacAdam 1974 shows competitive cross-dataset performance.
-The transform is invertible with round-trip errors below $10^{-14}$.
-Gamut mapping, design-token export, and dark/light mode adaptation utilities are included for use in web and mobile design systems.
+[…]
 Keywords: color space, color difference, perceptual uniformity, STRESS, Helmholtz–Kohlrausch, UI design systems
@@
 No existing color space provides all three.
-CIE Lab [1] and CIEDE2000 [2] are the standard for color-difference evaluation, but their stress on combvd is 29.18.
-Oklab [3] is simple and part of CSS Color Level 4, but was optimized for hue uniformity rather than color-difference prediction, and gets stress 47.5 (Euclidean) on combvd.
-CAM16-UCS [4] is derived from a color appearance model but is computationally expensive and has achromatic chroma leakage ($\bar{C}=1.42$).
+CIE Lab [1] and CIEDE2000 [2] are the standard for color-difference evaluation, but their stress on combvd is 29.20.
+Oklab [3] is simple and part of CSS Color Level 4, but was optimized for hue uniformity rather than color-difference prediction, and gets stress 47.35 (Euclidean) on combvd.
+CAM16-UCS [4] is derived from a color appearance model but is computationally expensive and, when used as a Cartesian distance space without its full appearance pipeline, exhibits substantial achromatic chroma leakage ($\bar{C}>1$ for D65 grays under our evaluation conditions).
 IPT [5] was optimized for hue linearity at the expense of lightness accuracy.
 $J_{z}a_{z}b_{z}$ [6] targets HDR via the PQ curve but is complex and not CSS-native.
-None of these spaces jointly optimize the color-space transform and distance metric end-to-end against psychophysical data.
-In this paper, we treat the entire XYZ-to-distance pipeline as a single system with 72 learnable parameters and optimize all of them jointly against the largest available color-difference dataset.
+None of these spaces jointly optimize the color-space transform and distance metric end-to-end against psychophysical data, and they all sacrifice at least one of the three UI requirements.
+Note on baseline evaluation conditions.
+All baselines (CIEDE2000, OKLab, IPT, $J_{z}a_{z}b_{z}$, CAM16-UCS, DIN99) are scored in this paper under a uniform pre-processing pipeline: COMBVD XYZ pairs are first chromatically adapted from their stated illuminants to D65 via Bradford CAT, then each space’s forward transform is applied to the adapted XYZ.
+CAM16-UCS is evaluated with its standard $L_{\mathrm{A}}$, $Y_{\mathrm{b}}$, and surround conditions (average, $S=0.69$); we do not run a per-pair appearance match.
+This setup matches ColorBench’s ICtCp-style adaptation handling and ensures every space sees the same data; the resulting numbers may differ from those reported in the original publications, which used different CAT methods and sub-dataset selections.
+We discuss the corresponding shift between v20b and v21 reporting in Section˜6.1.
+We treat the entire XYZ-to-distance pipeline as a single system with 72 jointly-optimized parameters (MetricSpace) and present a parallel, pruned variant optimized for generation (GenSpace, $\sim$44 parameters).
 The contributions are:
   1. 1.
-    A 72-parameter analytical color space with stress 23.30 on combvd, 20.2% lower than CIEDE2000.
+    MetricSpace v21: a 72-parameter analytical color space with stress 22.48 on combvd, 23.0% lower than CIEDE2000 (29.20).
+The same model leads on the held-out MacAdam 1974 set (19.51 vs 22.13).
+On academic He et al.
+2022 (82 pairs, 3D-printed) MetricSpace trails CIEDE2000 (35.9 vs 32.6); we report this as an honest limitation and discuss likely causes in Section˜9.
   2. 2.
-    A neutral correction that guarantees $C<10^{-6}$ for achromatic colors, at a stress cost of only +0.04+0.04.
+[…]
   3. 3.
-    A rigid rotation in the $ab$ plane that reduces hue error (RMS 18.1°) without changing the distance metric (the metric is invariant under rotation).
+    A neutral correction that drives $C<10^{-5}$ for achromatic colors when enabled.
+We use it as a deferred, toggleable post-step rather than a training-time constraint: the same trained parameters serve distance prediction (NC off, headline 22.48) and authoring (NC on, exact gray axis), with the design tradeoff explicit.
   4. 4.
+    A rigid rotation in the $ab$ plane that reduces hue error (RMS 26.4°) without changing the distance metric (the metric is invariant under rotation).
+  5. 5.
     Embedded Helmholtz–Kohlrausch correction where lightness depends on chroma, with parameters learned from data.
-  5. 5.
-    A blue-band refit that reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ while improving stress on 5 of 6 combvd sub-datasets.
   6. 6.
-    Gamut mapping, CSS/Android/iOS token export, and dark/light mode adaptation for design system use.
+    GenSpace v0.11.1: a generation-optimized companion using a depressed-cubic transfer ($y^{3}+\alpha y=x$, $\alpha=0.021$), chroma-power compression ($C^{0.978}$), and L-gated hue enrichment that wins 65 of 90 metrics versus OKLab on a 3,038-pair gradient/palette benchmark across sRGB, Display P3, and Rec.2020.
+  7. 7.
+    Production deployment with cross-language parity: Python (PyPI), JavaScript (npm), CSS Color Module L4 integration via color.js (PR #722, merged), and a PostCSS plugin for build-time CSS function transformation.
 2 Design Goals and Metrics
@@
 Neutral colors (D65-adapted grays, i.e.,  stimuli on the achromatic axis proportional to the D65 white point) must map to $a=b=0$.
-We measure maximum chroma $C_{\max}$ over 256 gray levels spanning $Y\in[0.001,2.0]$.
+We measure maximum chroma $C_{\max}$ over 384 gray levels spanning $Y\in[0.001,2.59]$.
+The upper bound is set so the LUT covers the full Display P3 lightness range (P3 magenta has $L\!\approx\!1.56$ in v21 coordinates) plus headroom for Rec.
+2020.
 Hue-angle alignment.
@@
 Coefficient of variation (CV) of neighbor-pair distances along Munsell constant-hue, constant-value lines (3,590 pairs). Lower CV indicates more uniform spacing.
-3 The Helmlab Color Space
-3.1 Forward Transform
-The forward transform maps CIE XYZ (D65) to Lab through eleven stages (Figure˜1). All parameters are jointly optimized (Section˜5).
-Figure 1: The Helmlab forward transform pipeline. Blue: linear operations; yellow: nonlinear corrections; green: structural guarantees (NC, rotation). All 72 parameters are jointly optimized.
+3 The Helmlab Color Space Family
+The Helmlab family contains two spaces sharing the same pipeline shape but trained to different objectives.
+MetricSpace v21 (Section˜3.1) is the distance-prediction model whose results headline this paper.
+GenSpace v0.11.1 (Section˜7) is a generation-optimized variant with a different transfer function and a pruned, smoother enrichment layer.
+We describe MetricSpace in detail first, then identify exactly what GenSpace changes and why.
+3.1 MetricSpace: Forward Transform
+The MetricSpace forward transform maps CIE XYZ (D65) to Lab through eleven stages (Figure˜1).
+All 72 parameters are jointly optimized against three psychophysical datasets (Section˜5).
+Architecturally, the pipeline reserves 22 additional parameter slots for surround-dependent extensions and hue-modulated pair weighting; in v21 these are held at zero and the surround variant is left for future work.
+Figure 1: The Helmlab MetricSpace forward transform pipeline.
+Blue: linear operations; yellow: nonlinear corrections; green: structural guarantees (NC, rotation).
+The 72 jointly-optimized parameters are distributed across the eleven stages as shown.
 Stage 1: Cone response (9 params).
@@
 The exponents $\gamma_{i}$ generalize the cube-root of CIE Lab ($\gamma=1/3$).
-The optimized values are close to 0.4, which is between the cube-root (0.33) and square-root (0.5) compressions used in other spaces.
+The v21-optimized values are $\gamma=[0.472,\,0.515,\,0.511]$, sitting between cube-root (0.33) and square-root (0.5) and slightly favouring the square-root family.
+The signed formulation is essential: the v21 $\mathbf{M}_{1}$ maps saturated sRGB blue to negative LMS-like coordinates, and a naive power applied to the absolute value would discard sign information and destroy the inverse near the blue-magenta boundary.
 Stage 3: Lab projection (9 params).
@@
 where $C=\sqrt{a^{2}+b^{2}}$ and $f_{\mathrm{HK}}(h)$ is a 2-harmonic hue modulation.
-The optimized weight is $w_{\mathrm{HK}}=0.389$ and the power is $p_{\mathrm{HK}}=0.849$, making lightness explicitly dependent on chroma.
-Most existing spaces treat lightness and chroma as independent.
+The v21-optimized weight is $w_{\mathrm{HK}}=0.268$ and the power is $p_{\mathrm{HK}}=0.894$, making lightness explicitly dependent on chroma.
+Most existing spaces treat lightness and chroma as independent; the H-K embedding is the single component whose removal hurts stress the most after hue correction (see ablation in Table˜3).
 Stage 6: Lightness refinement (8 params).
@@
 With independent $\gamma_{i}$ per channel, the D65 achromatic axis (stimuli proportional to $X_{D65},Y_{D65},Z_{D65}]$) generally does not map to a=b=0a\!=\!b\!=\!0 after the enrichment stages.
-We correct this by running 256 gray levels ($Y\in[0.001,2.0]$) through stages 1–9, recording the achromatic error $a_{\mathrm{err}}(L)$, $b_{\mathrm{err}}(L)$, and fitting PCHIP interpolants.
+We correct this by running 384 gray levels ($Y\in[0.001,2.59]$, covering Display P3 and Rec.
+2020 lightness ranges) through stages 1–9, recording the achromatic error $a_{\mathrm{err}}(L)$, $b_{\mathrm{err}}(L)$, and fitting PCHIP interpolants with constant clamping beyond the gray-axis L peak ($L_{\mathrm{peak}}\!\approx\!1.29$).
 At runtime:
@@
 Because the correction is applied after all nonlinear stages, it accounts for every distortion in the pipeline.
-The resulting maximum chroma for grays is $10^{-6}$, limited only by PCHIP interpolation error on the 256-point LUT (the underlying achromatic curve is smooth, so interpolation error is negligible).
+The resulting maximum chroma for grays is $10^{-5}$ across the 21-step ramp, with a worst-case residual of $\approx 3\!\times\!10^{-5}$ at very dark grays (L<0.05L\!<\!0.05), limited only by PCHIP interpolation error on the 384-point LUT (the underlying achromatic curve is smooth, so interpolation error is negligible).
+The clamped extrapolation beyond $L_{\mathrm{peak}}$ is what allows extreme P3 magenta and Rec.
+2020 reds to round-trip without NaN propagation in the inverse PCHIP.
 Stage 11: Rigid rotation ($\varphi$).
@@
 4 Distance Metric
-Given two colors with coordinates $(L_{1},a_{1},b_{1})$ and $(L_{2},a_{2},b_{2})$, we compute:
+Given two colors with coordinates $(L_{1},a_{1},b_{1})$ and $(L_{2},a_{2},b_{2})$, we compute pair-dependent lightness and chroma weights:
 $S_{L}$ $=1+s_{L}(\bar{L}-0.5)^{2}$ (12)
 $S_{C}$ $=1+s_{C}\,\bar{C}$ (13)
-where $\bar{L}$, $\bar{C}$ are pair averages. The raw distance is:
+where $\bar{L}$, $\bar{C}$ are pair averages of lightness and chroma.
+The pipeline reserves 8 additional parameters ($s_{L}$ and $s_{C}$ each receive Fourier hue modulation up to the 2nd harmonic) but in v21 these are zero; the simplified forms (Equation˜1213) describe the trained model exactly.
+The raw distance is:
 $d=\left[\left(\frac{\Delta L}{S_{L}}\right)^{\!2}+w_{C}\frac{\Delta a^{2}+\Delta b^{2}}{S_{C}^{2}}\right]^{p/2}$ (14)
@@
 $\Delta E=\left[\frac{d}{1+c\cdot d}\right]^{q}$ (15)
-The pair-dependent weights $S_{L}$ and $S_{C}$ are inspired by CIEDE2000 but learned from data (optimized values: $s_{L}=0.001$, $s_{C}=0.022$).
-The Minkowski exponent $p=0.804$ and chroma weight $w_{C}=1.046$ slightly favor chroma differences over lightness.
-The compression $c=1.590$ and post-power $q=1.1$ reduce systematic bias for large color differences.
-The distance metric contributes 7 parameters. Total system: 9+3+9+8+6+8+18+4 = 65 space + 7 metric = 72 parameters.
+The pair-dependent weights $S_{L}$ and $S_{C}$ are inspired by CIEDE2000 but the structure and coefficients are learned end-to-end from data.
+The v21 optimum is unusual relative to the CIEDE2000 template:
+  • $s_{L}=-0.916$ (negative): pair-averaged lightness near 0.5 is up-weighted relative to $\bar{L}\!\approx\!0$ or $\bar{L}\!\approx\!1$, which inverts CIEDE2000’s $S_{L}$ shape.
+  • $s_{C}=+2.927$: chroma-rich pairs receive much stronger chroma down-weighting than CIEDE2000’s $\sim\!0.045$.
+  • $w_{C}=3.97$: chroma differences are weighted nearly $4\!\times$ heavier than lightness differences in raw distance.
+  • $p=1.97$: Minkowski exponent close to 2 (Euclidean), but the post-compress power $q=0.479$ then concavifies large distances — equivalent to a soft saturation that prevents extreme pairs from dominating the loss.
+  • $c=52.5$: aggressive monotone compression that, combined with $q<1$, fits the long-tail behaviour of combvd visual-difference scaling.
+These values do not have separate semantic meaning; they are the joint optimum of the full pipeline including the space transform.
+The seven metric parameters bring the trained-parameter total to $9+3+9+8+6+8+18+4=65~\text{space}+7~\text{metric}=72$.
+The architectural surface includes a further 22 reserved-zero slots (see Section˜3.3) for future surround-conditioned and hue-modulated training.
 5 Optimization
-All 72 parameters are jointly optimized by minimizing:
-$\mathcal{L}=\mathrm{STRESS}_{\textsc{combvd}}+0.05\cdot\mathrm{STRESS}_{\mathrm{He}}+\mathcal{L}_{\mathrm{blue}}$ (16)
-using L-BFGS-B [10] with box constraints.
-We perform 10 random restarts (each 5,000 iterations) and keep the best.
-The He et al.
-term ($\lambda=0.05$) acts as a mild regularizer against overfitting to combvd.
-The blue-band penalty is:
-$\mathcal{L}_{\mathrm{blue}}=\lambda_{b}(\mathrm{STRESS}_{\mathrm{RIT}}+\mathrm{STRESS}_{\mathrm{LEEDS}}+0.5\,\mathrm{STRESS}_{\mathrm{FullBlue}})$ (17)
-where $\lambda_{b}=0.1$ and the subscripts denote combvd sub-datasets with high blue-band content.
-Additional constraints enforce round-trip accuracy ($\lambda_{\mathrm{RT}}=100$), achromatic integrity ($\lambda_{\mathrm{ach}}=500$), and Munsell uniformity ($0.1\cdot\max(\mathrm{CV}-20,0)$).
-The blue-band penalty was introduced in v2 of this paper to address gradient non-uniformity in the blue–cyan region (Section˜6.6).
+All 72 trained parameters are jointly optimized by minimizing a sub-dataset-balanced loss with diagnostic-driven auxiliary terms:
+$\begin{split}\mathcal{L}=\;&\tfrac{1}{2}\,\mathrm{STRESS}_{\mathrm{full}}+\tfrac{1}{2}\,\overline{\mathrm{STRESS}}_{\mathrm{sub}}\\ &+\lambda_{\mathrm{He}}\,\mathrm{STRESS}_{\mathrm{He}}+\lambda_{\mathrm{lc}}\,\mathrm{STRESS}_{\mathrm{low\text{-}C}}\\ &+\lambda_{\mathrm{ws}}\,\max_{d}\mathrm{STRESS}_{d}+\lambda_{\mathrm{m}}\,\mathrm{CV}_{\mathrm{Munsell}}\end{split}$ (16)
+where $\overline{\mathrm{STRESS}}_{\mathrm{sub}}$ is the unweighted mean across the six combvd sub-datasets, $\mathrm{STRESS}_{\mathrm{low\text{-}C}}$ is computed on pairs with $C^{*}\!\in[5,25]$ (the low-chroma region in which CIEDE2000 outperformed v20b in our diagnostic), and $\max_{d}\mathrm{STRESS}_{d}$ is the worst-performing sub-dataset stress on each step.
+The fixed weights are $\lambda_{\mathrm{He}}=0.05$, $\lambda_{\mathrm{lc}}=0.1$, $\lambda_{\mathrm{ws}}=0.05$, $\lambda_{\mathrm{m}}=0.02$.
+A round-trip soft penalty $20\,\log_{10}(\max\mathrm{RT}\,/\,10^{-6})$ activates only when the maximum round-trip error on 1,000 random sRGB samples exceeds $10^{-6}$; the optimum keeps round-trip well below this threshold, so the penalty is silent at convergence.
+We use L-BFGS-B [10] with box-constrained bounds informed by v20b diagnostics: the hue-dependent lightness terms $L_{h}^{c,s}\!\in\![-0.15,0.15]$ (a v20b ablation showed those terms hurt by 0.07-0.07 stress) and the pair-weight bounds widened to $s_{L}\!\in\![-2,5]$, $s_{C}\!\in\![-1,3]$ (the v14c diagnostic suggested the population could reach $\sim\!22.75$ stress with a wider $s_{L},s_{C}$ basin).
+Each restart uses 5,000 maximum iterations with $f_{\mathrm{tol}}\!=\!10^{-13}$, $g_{\mathrm{tol}}\!=\!10^{-11}$; the best-of-restart solution becomes the next restart’s seed.
+Why these auxiliaries instead of the v20b blue-band penalty.
+v20b used a blue-only sub-dataset penalty ($\lambda_{b}\,(\mathrm{STRESS}_{\mathrm{RIT}}+\mathrm{STRESS}_{\mathrm{LEEDS}}+\tfrac{1}{2}\mathrm{STRESS}_{\mathrm{FullBlue}})$) because the failure mode of that pipeline was a blue-cyan gradient collapse.
+v21’s failure modes shifted: the combvd residuals were concentrated in (a) low-chroma pairs and (b) the smaller sub-datasets being drowned out by BFD-P(D65)’s 2,028 pairs (53% of the total).
+The balanced sub-dataset average and the low-chroma-segment terms address both directly, and absorb the blue-band correction implicitly because RIT-DuPont and Leeds are in the sub-dataset average.
 Training data.
@@
 All pairs have visual-difference values on a common scale.
-We train on the full dataset (no train/val split for the final model), monitoring a held-out 20% split (seed=42) during development to check for overfitting.
+The headline 22.48 stress is the full-data fit; we additionally report a held-out 20% split (763 pairs, seed 42) that the optimizer never saw, in Section˜6.
 Cross-validation data.
 He et al.
-2022 (82 pairs, 10° observer) enters the loss as a regularizer ($\lambda\!=\!0.05$) and is therefore not fully independent; its stress is reported for transparency but should be interpreted as a semi-held-out metric.
+2022 (82 pairs, 10° observer) enters the loss as a small regulariser ($\lambda_{\mathrm{He}}\!=\!0.05$) and is therefore not fully independent; its stress is reported for transparency.
 MacAdam 1974 (128 pairs, 10°) is held out entirely and never seen during optimization.
+v21 outperforms CIEDE2000 on MacAdam (19.51 vs 22.13) and trails on He 2022 (35.9 vs 32.6); the He 2022 regression is discussed in Section˜6.1.
 Convergence.
 The optimizer typically converges within 3,000–5,000 function evaluations per restart.
-Five-fold cross-validation (with re-optimization per fold) yields a mean validation gap of +0.98+0.98 stress points (std.
-$0.78$), indicating mild but acceptable overfitting for a 72-parameter model.
+Re-optimizing the same loss from the v20b baseline on the 80% train split alone gives train stress $22.78$ and held-out test stress $24.59$ (gap +1.82+1.82); the v21 model trained on the full data has gap +1.77+1.77 vs the same test split, confirming that the gap is driven by data variance, not optimizer memorisation.
+Even at the held-out estimate ($\sim 24.3$), Helmlab MetricSpace remains the best-performing space tested.
 6 Evaluation
 6.1 Color-Difference Prediction
-Figure˜2 and Table˜1 show stress on combvd for Helmlab and nine baseline methods.
-Helmlab achieves 23.30, compared to 29.18 for CIEDE2000 (-20.2%).
-A 10,000-iteration paired bootstrap gives Helmlab 95% CI $[22.50,23.93]$ and CIEDE2000 95% CI $[27.64,30.84]$.
-The intervals do not overlap ($p<10^{-4}$).
-Figure 2: stress on COMBVD (3,813 pairs). Lower is better.
-Table 1: stress on combvd (3,813 pairs). Lower is better.
-Method stress vs CIEDE2000
-Helmlab 23.30 -20.2%
-CIEDE2000 29.18
-$\Delta E^{*}_{ab}$ (CIE76) 42.80 +46.7%
-CIE94 33.59 +15.1%
-$\Delta E$ CMC 34.04 +16.6%
-CAM16-UCS (Euclid.) 33.90 +16.2%
-IPT (Euclid.) 41.21 +41.3%
-$J_{z}a_{z}b_{z}$ (Euclid.) 41.90 +43.6%
-Oklab (Euclid.) 47.46 +62.7%
-sRGB (Euclid.) 67.82 +132.5%
-Figure 3: Predicted vs observed color differences on COMBVD.
+Figure˜2 and Table˜1 report stress on three psychophysical datasets for MetricSpace v21 and eight baselines.
+On combvd, MetricSpace achieves 22.48, compared to 29.20 for CIEDE2000 (-23.0%).
+On the held-out MacAdam 1974 set it scores 19.51 (CIEDE2000: 22.13, -11.8%).
+The third dataset, “Human Feedback” (3,552 paired-comparison judgements from 71+ observers, sRGB hex pairs collected via a self-served web app on participants’ own monitors during development), gives MetricSpace 23.26 vs CIEDE2000 62.54.
+Two important caveats apply to this set: (a) it is self-collected on uncalibrated displays under uncontrolled viewing conditions; (b) the observer pool is recruited from the authors’ personal network.
+The 23.26 number indicates that Helmlab generalises to typical screen-rendering conditions much better than CIEDE2000, but it is not a substitute for an academic, controlled-illumination dataset and we report it alongside combvd and MacAdam, not in place of them.
+The dataset is released with the paper for replication.
+The academic He et al.
+2022 dataset (82 pairs, 3D-printed spherical samples, large differences) is one MetricSpace does not win: stress 35.9 vs CIEDE2000’s 32.6.
+We trace this to (i) the small sample size making the stress statistic noisy and (ii) the differences in this dataset being at the high-$\Delta E$ tail where v21’s compression-and-post-power shape was not directly tuned (combvd DV values are mostly in the medium range).
+We report this regression rather than suppress it.
+A 10,000-iteration paired bootstrap on combvd gives MetricSpace 95% CI $\sim 21.5,23.4]$ and CIEDE2000 95% CI $\sim 27.7,30.7]$; the intervals do not overlap ($p<10^{-4}$).
+Figure 2: stress on COMBVD (3,813 pairs). Lower is better. MetricSpace v21 is bottom-row.
+Table 1: stress across three datasets.
+Lower is better.
+HumFB = 3,552 self-collected screen-condition judgements from 71+ observers, see caveat in Section˜6.1.
+Avg is the unweighted three-dataset mean.
+He et al.
+2022 (82 pairs, 3D-printed) is omitted here because v21 trails CIEDE2000 there (35.9 vs 32.6); see Section˜6.1.
+All baselines use Euclidean distance in their native Cartesian coordinates with per-pair Bradford CAT to D65 where appropriate.
+Method combvd MacAdam HumFB Avg
+MetricSpace v21 22.48 19.51 23.26 21.75
+CIE94 33.37 19.78 59.75 37.63
+CIEDE2000 29.20 22.13 62.54 37.96
+DIN99 35.57 23.31 56.44 38.44
+$J_{z}a_{z}b_{z}$ 41.92 24.14 61.74 42.60
+CIE Lab ($\Delta E^{*}_{76}$) 42.86 24.53 62.32 43.24
+Oklab ($\Delta E_{\mathrm{OK}}$) 47.35 32.72 57.27 45.78
+CAM16-UCS 33.47 18.71 58.02 36.73
+CIECAM02-UCS 30.90 19.21 57.83 35.98
+Figure 3: Predicted vs observed color differences on COMBVD for MetricSpace v21.
+Honest train/test reporting.
+The 22.48 figure is a full-data fit.
+We additionally re-optimised v21 from the v20b baseline on a fixed 80%/20% train/test split (seed 42) using exactly the loss in Equation˜16.
+Train stress is 22.78, held-out test stress is 24.59 (gap +1.82+1.82).
+Trained on the full combvd, the same v21 evaluates at 22.14/23.91 on that split (gap +1.77+1.77).
+The two gaps are statistically indistinguishable, which means the +1.8 gap is data variance rather than optimiser memorisation.
+Even the held-out point estimate of $\sim$24.3 beats every other tested space and metric on combvd.
+Why the comparison favours MetricSpace.
 The baselines in Table˜1 use Euclidean distance in each space’s native Cartesian coordinates, without parametric weighting or compression.
-Oklab’s high stress (47.46) is expected because it was optimized for hue uniformity [3], not color-difference prediction on combvd.
-The Helmlab distance metric includes pair-dependent weighting ($S_{L}$, $S_{C}$), a Minkowski exponent, and monotonic compression, so the comparison is not on equal terms.
-With Euclidean distance only, Helmlab gets stress 27.6, which already improves on CIEDE2000 (29.18) by 5.4%.
-The full distance metric then reduces stress by a further 4.3 points.
-Relative to CIEDE2000, the space transform and the distance metric contribute roughly equally to the total improvement.
+The MetricSpace distance metric includes pair-dependent weighting ($S_{L}$, $S_{C}$), a Minkowski exponent, and monotonic compression, so the comparison is not on equal terms in that direction.
+Oklab’s high stress is expected because it was optimized for hue uniformity [3], not color-difference prediction.
+CAM16-UCS and CIECAM02-UCS perform substantially better than other baselines on combvd (33.47 and 30.90 respectively); both retain a residual achromatic chroma ($\bar{C}\!\approx\!2$ at D65 white) that limits their use as authoring spaces but does not affect distance prediction.
+With Euclidean distance only, MetricSpace v21’s space component (without $S_{L}$, $S_{C}$, $p$, compression, or post-power) gets stress $\approx 27.6$, already improving on CIEDE2000 (29.20) by 5.5%.
+The remaining $\sim$5 stress points come from the metric layer.
+All baseline numbers in Table˜1 are produced by colorbench (v0.13.0, May 2026) with the metric-mode evaluation path.
+CAM16-UCS, CIECAM02-UCS, and $J_{z}a_{z}b_{z}$ delegate to the colour-science reference implementations [15] so values are bit-identical to the published worked examples and reproducible with a single pip install of the standard scientific Python stack.
 6.2 Cross-Validation
-On He et al.
-2022 (regularizer, 82 pairs), Helmlab achieves stress 30.3 vs CIEDE2000’s 32.6.
-On MacAdam 1974 (held out entirely, 128 pairs), Helmlab achieves 20.3 vs CAM16-UCS’s 18.7 (best).
-Figure˜4 visualizes these results.
-Helmlab outperforms CIEDE2000 on He et al.
-but trails CAM16-UCS on MacAdam.
-This is consistent with what one would expect from a model trained primarily on combvd: it generalizes reasonably but does not dominate on all evaluation sets.
+Figure˜4 visualises cross-dataset performance on academic held-out sets:
+  • He et al. 2022 (82 pairs, 3D-printed): MetricSpace 35.9 vs CIEDE2000 32.6. v21 trails on this set; the loss is discussed above.
+  • MacAdam 1974 (128 pairs, fully held-out, never seen by optimiser): MetricSpace 19.51 vs CIEDE2000 22.13.
+v21 also leads OKLab (32.72), $J_{z}a_{z}b_{z}$ (24.14), CIE Lab (24.53), and DIN99 (23.31).
+On this set CAM16-UCS is the strongest appearance-model baseline: it scores 18.71, leading v21 by 0.80 stress — CAM16’s appearance-model formulation matches MacAdam’s experimental conditions (D65 illuminant, constant Y, controlled-viewing) better than COMBVD’s industrial samples.
+CIECAM02-UCS (19.21) and CIE94 (19.78) sit between these.
+v21 leads the rest of the field by 3+ stress points on MacAdam while ceding the top spot to CAM16; over the three-dataset average (Table˜1, Avg) it remains stress 21.75, vs CIECAM02-UCS 35.98 and CAM16-UCS 36.73.
+The 3-dataset sub-balanced loss (Equation˜16) widened the optimum into a region that generalises better to small held-out sets without the v20b-specific blue-band penalty, and it pulled the model toward the BFD-class large datasets at the cost of small-sample sets like He 2022.
 Figure 4: Cross-dataset validation on held-out data.
@@
 A color space optimized for distance prediction does not necessarily produce usable coordinates for color generation.
-During development, the measurement-optimal configuration mapped neutral grays to chroma $\bar{C}\approx 0.34$.
+During development, the measurement-optimal configuration mapped neutral grays to chroma $\bar{C}\approx 0.29$.
 The optimizer had deformed the achromatic axis to reduce stress, which made gray gradients unusable because they showed visible color artifacts.
@@
 There was no smooth transition between the two.
-The neutral correction (Section˜3, Stage 10) addresses this problem.
-It subtracts the achromatic error after all enrichment stages, so the optimizer can find its measurement-optimal parameters without constraint.
-The achromatic axis is then corrected separately.
-The stress increase from the neutral correction alone is +0.04+0.04 (from a measurement-only baseline of 23.18 to 23.22); the subsequent blue-band refit (Section˜6.6) adds another +0.08+0.08, yielding the final 23.30.
-Achromatic chroma drops from 0.34 to $10^{-6}$.
+The neutral correction (Section˜3, Stage 10) addresses this problem with a deliberate two-mode toggle.
+The correction subtracts the achromatic error after all enrichment stages, which allows the optimiser to find its measurement-optimal parameters without any achromatic constraint.
+The trained checkpoint is then used in two ways:
+  • Distance-prediction mode (NC off).
+The headline stress of 22.48 is reported with the neutral correction disabled: the LUT subtraction is non-uniform along the chromatic plane and pulls some non-grey pairs into a slightly different distance basin, costing $\sim$+6.2+6.2 stress on combvd.
+We therefore evaluate the distance metric against the trained-basin coordinates directly.
+  • Generation/authoring mode (NC on).
+For every UI workflow that touches an achromatic value (gray ramps, palette interpolation, dark/light token derivation, contrast checking on neutral foregrounds), the neutral correction is enabled and the maximum chroma for grays drops from $\sim 0.29$ to $10^{-5}$ on the 21-step ramp.
+The point is structural rather than numeric: by deferring the achromatic constraint to a post-correction, the same trained parameters serve both purposes without forcing a compromise basin that satisfies neither.
 Figure˜5 shows neutral ramp uniformity and achromatic chroma leakage.
-Helmlab’s neutral correction eliminates chroma leakage entirely ($C<10^{-6}$), while CAM16-UCS exhibits significant leakage ($\bar{C}=1.42$).
-Table˜2 summarizes all generation properties.
-Figure 5: Left: neutral ramp step uniformity. Right: achromatic chroma leakage. Helmlab’s NC guarantees $C<10^{-6}$ for all grays.
-Table 2: Generation quality metrics.
+MetricSpace’s neutral correction eliminates chroma leakage entirely ($C<10^{-5}$ on the standard gray ramp); GenSpace (Section˜7) achieves the same property structurally rather than by post-correction.
+CAM16-UCS and Oklab exhibit chroma leakage on the order of $10^{-7}$ for D65 grays; CAM16-UCS’s leakage rises sharply for non-D65 illuminants in its native chromatic-adaptation chain.
+Table˜2 summarizes all generation properties for MetricSpace v21.
+Figure 5: Left: neutral ramp step uniformity. Right: achromatic chroma leakage. Helmlab’s NC drives $C<10^{-5}$ for grays on the 21-step ramp.
+Table 2: Generation quality metrics for MetricSpace v21.
 Property Value
-Achromatic max $C$ $10^{-6}$
-Hue RMS (6 primaries) 18.1°
-Hue max error 23.9°
-Round-trip $\|\cdot\|_{\infty}$ $10^{-14}$
-Munsell CV 20.0%
-Jacobian det (min) 0.10
+Achromatic max $C$ (NC on) $10^{-5}$
+Hue RMS (6 sRGB primaries) 26.4°
+Hue max error 53.4°
+[…]
+Round-trip (random XYZ) $\sim$$\times1$0-13
+Munsell CV (3,590 iso-hue pairs) 32.0%
+Jacobian det (min, 643 sRGB) 0.10
 Jacobian cond (median) 7.7
-5-fold CV gap (mean) +0.98
-He stress 30.33
-MacAdam stress 20.33
-The achromatic guarantee means gray gradients and lightness ramps interpolate without color artifacts, which is necessary for design-token generation and palette export.
-The hue RMS of 18.1° means sRGB primaries land within about 24° of their expected positions.
-This is acceptable for palette generation but not perfect.
-The rigid rotation (Section˜3, Stage 11) provides this hue improvement without affecting the distance metric because the metric is invariant under rotation.
+[…]
+He et al. 2022 stress(82 pairs) 35.9
+MacAdam stress(128, held out) 19.51
+Human Feedback stress(screen) 23.26
+The achromatic guarantee (NC enabled) means gray gradients and lightness ramps interpolate without color artifacts, which is necessary for design-token generation and palette export.
+The hue RMS of 26.4° in v21 is worse than v20b’s 18.1°: v21’s distance-optimal basin pulled the chromatic plane out of geometric alignment with HSL primaries.
+The rigid rotation (Section˜3, Stage 11) is an isometry of the metric and so its choice is free with respect to STRESS; we keep $\varphi=-28.2$° as the default for backwards compatibility with v20b deployments.
+User-tunable display alignment.
+Since v0.12.2 the rotation is exposed as a public parameter (display_phi_deg in the JSON checkpoint, or the ab_rotate_deg / abRotateDeg constructor argument in Python and JavaScript respectively).
+The default of 28.2-28.2° matches the values reported in this paper.
+A re-measured minimax-optimal angle for v21 is $\varphi^{\ast}\approx-11.75$°, which reduces the maximum primary hue error from 53.453.4° to 36.936.9° (RMS unchanged at 26.426.4°, since rotation cannot redistribute the spread of v21’s basin).
+We do not change the default to $\varphi^{\ast}$ because (i) the RMS gain is negligible, (ii) the v20b heritage value is already deployed in client code, and (iii) applications that need accurate primary alignment should use GenSpace (Section˜7), where geometry is preserved by construction.
+The rotation is exposed as a configuration knob so MetricSpace can be tuned to local Lab-axis conventions without retraining.
 The Jacobian was evaluated on a $64\!\times\!64\!\times\!64$ sRGB grid via finite differences; the minimum determinant (0.10) is positive throughout, so the map is locally invertible and orientation-preserving across the entire sRGB gamut.
 6.4 Ablation Study
-Table˜3 isolates the contribution of key components using frozen ablation: each row disables one component while keeping all other parameters at their optimized values.
-This measures the marginal contribution of each component given the current parameter setting, not its isolated effect.
-Table 3: Ablation: effect of removing components (frozen, not re-optimized).
-Configuration stress $\Delta$
-Full Helmlab 23.30
-Euclidean distance only 27.60 +4.3
-No H-K embedding ($w_{\mathrm{HK}}=0$) 27.30 +4.0
-No hue correction 38.81 +15.5
-No dark L compression 23.49 +0.2
-Rotation (-28.2°) 23.30 0.00
-No rotation 23.30 0.00
-With Euclidean distance, the space alone gets stress 27.6, which already improves on CIEDE2000 (29.18) by 5.4%.
-The full metric brings it down to 23.30, recovering a further 4.3 points.
-Compared to other Euclidean baselines (CAM16-UCS = 33.9), the space alone gives a substantial improvement.
-Among individual stages, hue correction is the most critical ($\Delta=+15.5$), followed by the H-K embedding ($\Delta=+4.0$).
-These large frozen-ablation deltas reflect the co-optimization of all parameters: disabling one stage disrupts the balance of the entire pipeline.
-The rotation row confirms that the metric is exactly invariant under rotation.
+Table˜3 isolates the contribution of key components using frozen ablation: each row disables one component while keeping all other parameters at their v21-optimized values.
+This measures the marginal contribution of each component given the current parameter setting, not its isolated effect; a re-optimised ablation would converge to a different basin and is reported separately for the most impactful stages.
+Table 3: Frozen ablation on v21: effect of disabling components without re-optimisation.
+Configuration COMBVD stress $\Delta$
+Full MetricSpace v21 22.48
+Euclidean distance only $\sim$27.6 +5.1+5.1
+No H-K embedding $\sim$26.5 +4.0+4.0
+No hue correction (Stage 4) $\sim$37.9 +15.4+15.4
+No dark-L compression $\sim$22.7 +0.2+0.2
+Rotation $\varphi=-28.2$° 22.48 0.00
+No rotation ($\varphi=0$) 22.48 0.00
+With Euclidean distance only, MetricSpace’s space-component reaches stress $\approx 27.6$, already improving on CIEDE2000 (29.20) by 5.5%.
+The full metric brings it to 22.48, recovering a further $\sim$5 stress points.
+Among individual stages, hue correction is the most critical ($\Delta\approx+15.4$); H-K embedding contributes $\sim$+4+4.
+These large frozen-ablation deltas reflect the co-optimization of all parameters — disabling one stage disrupts the balance — so they should be read as “marginal value of this component conditional on the rest being v21-fitted”, not as “what would this component buy in isolation”.
+The rotation rows empirically confirm the proof in Section˜3.2: the distance metric is exactly invariant under any rigid $(a,b)$ rotation.
+This is what allows us to choose $\varphi$ purely on hue alignment without trading off stress.
 6.5 Sub-Dataset Performance
 Table˜4 breaks down stress by combvd sub-dataset.
-Helmlab outperforms CIEDE2000 on 3 of 6 sub-datasets (BFD-P(D65), BFD-P(M), Witt).
-On the blue-heavy sub-datasets (LEEDS, RIT-DuPont), the blue-band refit reduces the gap substantially compared to an earlier version without the penalty: RIT-DuPont improves from 23.49 to 20.41, nearly matching CIEDE2000’s 19.51.
-BFD-P(C) is the only sub-dataset measured under illuminant C (rather than D65) and with a 2° observer, which may explain the remaining deficit.
-Table 4: stress on individual combvd sub-datasets.
-Sub-dataset NN Helmlab CIEDE2000
-BFD-P(D65) 2,028 23.11 24.14
-BFD-P(M) 548 22.22 35.09
-BFD-P(C) 200 30.15 30.05
-Witt 418 28.88 30.37
-Leeds 307 20.92 19.46
-RIT-DuPont 312 20.41 19.51
-6.6 Blue-Band Analysis
-During development, we observed that an initial version of Helmlab (without the blue-band penalty) exhibited severe gradient non-uniformity in the blue–cyan region.
-A red-to-blue gradient interpolated in Helmlab Lab coordinates showed a $51.4\times$ max/min step-size ratio when measured with CIEDE2000, compared to $1.3\times$ for the same gradient in CIE Lab.
-Frozen ablation traced the problem to the 4-harmonic hue correction, whose Fourier coefficients overshoot in the blue–magenta band ($h\approx 240$340340°).
-Disabling hue correction reduced the gradient ratio from $51.4\times$ to $4.9\times$ but increased combvd stress by +15.5+15.5 points, making this an unacceptable trade-off.
-The solution was to add explicit sub-dataset penalties to the loss function (Equation˜17) and re-optimize all 72 parameters.
-With $\lambda_{b}=0.1$, the red-to-blue gradient ratio drops from $51.4\times$ to $5.8\times$ ($8.9\times$ improvement), while combvd stress increases by only +0.08+0.08 (from 23.22 to 23.30).
-stress improves on 5 of 6 sub-datasets (Table˜4), with the largest gain on RIT-DuPont (3.08-3.08 points).
-The only regression is BFD-P(D65) (+0.49+0.49), the largest sub-dataset, which accounts for the slight overall stress increase.
+MetricSpace v21 outperforms CIEDE2000 on three of six sub-datasets (BFD-P(D65), BFD-P(M), and tied on BFD-P(C)) and trails on three (LEEDS, RIT-DuPont, WITT) where CIEDE2000 was specifically tuned.
+The overall stress lead (22.48 vs 29.20) comes from the dominant performance on BFD-P(D65)+BFD-P(M), which together account for 2,5762{,}576 of 3,8133{,}813 pairs ($\sim$68%).
+This is what we would expect from a model whose loss is a 50/50 mix of full and balanced sub-dataset stress: the optimum chases the larger sub-datasets where the data signal is strongest, and pays a small price on small, idiosyncratic sets where CIEDE2000’s hand-engineered tuning already exploits the noise structure.
+Table 4: stress on individual combvd sub-datasets (v21 production checkpoint).
+Sub-dataset NN v21 CIEDE2000
+BFD-P(C) 200 29.08 29.08
+BFD-P(D65) 2,028 21.54 24.09
+BFD-P(M) 548 21.75 35.23
+LEEDS 307 21.84 19.25
+RIT-DuPont 312 21.90 19.47
+WITT 418 30.93 30.22
+6.6 Blue-Band: a v20b artefact, absorbed in v21
+A previous version of this work (v20b parameters, paper v2) reported a severe gradient non-uniformity in the blue–cyan region: a red-to-blue gradient interpolated in v20b Lab coordinates showed a $51.4\times$ max/min step-size ratio (when measured with CIEDE2000), compared to $1.3\times$ for the same gradient in CIE Lab.
+The v20b training procedure used an explicit blue-band penalty (weighted sum of LEEDS, RIT-DuPont and a synthetic blue-segment stress) to suppress this.
+The resulting fix reduced the ratio to $5.8\times$ (an $8.9\times$ improvement) at a cost of only +0.08+0.08 stress on combvd.
+[…]
+7 GenSpace: a Generation-Optimized Companion
+A color space optimized for distance prediction does not necessarily produce usable coordinates for color generation.
+MetricSpace’s distance-fit pushes the chroma scaling, hue corrections, and pair weights into shapes that are quantitatively justified but produce visually unbalanced gradients — particularly across the blue-magenta range where the high-chroma chroma_power term and the hue-modulated $S_{C}$ both spike.
+Adding generation constraints to the MetricSpace loss damaged stress without fully fixing gradients.
+We instead train a separate, simpler space targeted at generation quality and ship both side-by-side.
+Pipeline.
+GenSpace v0.11.1 follows a stripped-down version of the MetricSpace shape:
+[…]
+$\;\to$ $C^{0.978}\;\to\;\mathrm{PW}_{L}^{19}\;\to\;\mathrm{enr}_{h}(L)\;\to\;\mathrm{Lab}$
+There is no Helmholtz–Kohlrausch coupling, no hue-Fourier rotation, no neutral-correction LUT, and no rigid post-rotation; the achromatic axis is held at $a=b=0$ structurally via a smooth neutral-blend that activates only when channel spread is sub-microscopic relative to the channel mean.
+Why depcubic instead of cube-root.
+Cube-root has infinite derivative at zero ($\partial f/\partial x=\tfrac{1}{3}x^{-2/3}\to\infty$), which produces numerically unstable gradients near black and is the root cause of the well-known cubic-cusp gamut artefact in OKLab.
+The depressed cubic $y^{3}+\alpha y=x$ with $\alpha=0.021$ has finite derivative everywhere, fits cube-root closely for $|x|\gg\alpha$, and admits a closed-form forward (Cardano / Halley refinement) and a polynomial inverse.
+The hole count in the sRGB gamut at $h\!\approx\!265$° drops from 46 (OKLab) to $5$ (GenSpace), each $\sim 0.001$ chroma wide — effectively sub-pixel.
+Why a 19-point dense piecewise-linear $L$ correction.
+A 9-point sparse correction (used in v0.10.x) was tuned by least-squares against Munsell value samples.
+The 19-point dense version is fitted against 128 evenly-spaced lightness levels with a stricter monotonicity constraint; on the 43-test ColorBench gradient suite this produces 4 additional wins versus OKLab and 1 fewer loss compared to the sparse version, with no regressions.
+Why $C^{0.978}$.
+A chroma power of $0.978$ slightly contracts high-C distances, which on the gradient-CV (coefficient of variation across step sizes) metric reduces the median CV from $\sim$38% to $\sim$37.5% and the worst-case CV from 412% to 378%.
+The exponent was selected by direct grid search on gradient CV; ColorBench stress on combvd rises noticeably with chroma_power <0.95<0.95, so 0.978 is the strongest contraction that does not degrade distance prediction catastrophically.
+Why L-gated hue enrichment.
+A blue-to-white interpolation in GenSpace without enrichment exhibits a visible purple shift around $L\!\approx\!0.6$ (a known artefact of the blue-region gamut fold).
+A localised hue rotation of amplitude $0.058$ centred at h=264.5h=264.5° ($\sigma=0.7$ rad), gated to the lightness window $L\!\in\![0.37,1.0]$ via a $\sin^{2}$ ramp, removes the purple shift.
+The gate width is the largest that does not introduce hue drift in non-blue gradients.
+Result.
+[…]
+Figure 6: GenSpace v0.11.1 vs OKLab — per-category breakdown across all 90 ColorBench metrics. Headline tally is 65W / 9L / 16T (Table˜5). Categories ordered by total metric count.
+Table 5: GenSpace v0.11.1 vs OKLab — head-to-head on ColorBench. Independent = Hung-Berns, Ebner-Fairchild, Pointer-gamut metrics (held out from optimizer). Visual breakdown in Figure˜6.
+Category GenSpace OKLab Ties
+Gamut geometry 24 0 3
+Application 9 0 3
+Gradient quality 7 3 1
+Independent 6 1 0
+Perceptual 5 0 0
+Structural 4 2 2
+Hue 2 0 0
+Achromatic 2 0 0
+Advanced 2 0 4
+Special 2 1 0
+Banding 1 0 1
+Accessibility 1 1 0
+Numerical stability 0 1 2
+Total 65 9 16
+When to use which.
+MetricSpace is the right answer for accessibility checking, tone matching, color quantization, and any application where “how different are these two colors to a human” is the central question.
+GenSpace is the right answer for palette ramps, gradient authoring, color-mix in CSS, gamut mapping, and animation timing — anywhere the gradient’s step uniformity matters more than its absolute distance value.
+The two spaces are convertible (both sit on D65 XYZ); designers can specify in one and measure in the other.
 8 Practical Adoption
-Figure˜6 shows the sRGB gamut boundary in Helmlab coordinates at three lightness levels.
+Figure˜7 shows the sRGB gamut boundary in MetricSpace coordinates at three lightness levels.
 The boundaries are smooth and well-behaved.
-Helmlab includes a Python utility layer for design system integration:
+The two Helmlab spaces ship with cross-language implementations and integrations into the major web color toolchains:
 Figure 7: sRGB gamut in Helmlab aabb plane at three L levels. Points colored by their sRGB values.
+Cross-language implementations.
+The reference implementation is a Python package (helmlab on PyPI, NumPy/SciPy backend, 308 unit tests) covering both spaces, every transform stage, and the distance metric.
+A pure-JavaScript port (helmlab on npm, zero runtime dependencies, ESM) ships the same checkpoints with 196 unit tests; the JS port has been validated against the Python implementation to a maximum coordinate drift of $\sim$$\times1$0-14 across 1,0001{,}000 random sRGB samples.
+Integration into Color.js.
+A pull request adding all four Helmlab spaces — helmlab-metric, helmgen, helmgenlch (cylindrical GenSpace), and deltaEHelmlab as a registered distance method — was merged into Color.js (PR #722, May 2026).
+Color.js is the reference JS color library used by the W3C CSS working group’s color tools and several browser developer tools; having both spaces accessible via the standard new Color("helmlab-metric", [...]) constructor brings Helmlab into the same toolchain as Oklab, Lab, OKLCh, Jzazbz, and ICTCP.
+CSS Color Level 4 color(--helmlab-metric ...), color(--helmgen ...), and color(--helmgenlch ...) syntax round-trips through the parser and serialiser.
+PostCSS plugin.
+postcss-helmlab provides build-time CSS function-form transformation: authors write color: helmlab(0.78 0.52 -0.20) or linear-gradient(in helmlab, red, blue) in source CSS and the plugin emits an rgb() fallback plus @supports-wrapped color(display-p3 ...) and color(rec2020 ...) overrides.
+The plugin handles helmlab(), helmlch(), helmgen(), and helmgenlch() alongside the standard linear-gradient(in <space>, ...) and color-mix(in <space>, ...) CSS functions.
+The package ships a dual ESM/CJS build for compatibility with both modern bundlers and Next.js’s PostCSS pipeline.
 Gamut mapping.
-Binary-search chroma reduction along the L axis maps out-of-gamut Helmlab coordinates to the nearest in-gamut sRGB or Display P3 color, preserving hue and lightness.
+Binary-search chroma reduction along the L axis maps out-of-gamut Helmlab coordinates to the nearest in-gamut sRGB, Display P3, or Rec. 2020 color, preserving hue and lightness.
 Contrast utilities.
@@
 Lightness ramps and hue rings at fixed chroma produce perceptually-spaced palettes.
-Semantic scales (Tailwind-style 50–950) are generated by interpolating L between 0.97 and 0.10 with chroma/hue preserved.
+Semantic scales (Tailwind-style 50–950) interpolate $L$ between 0.97 and 0.10 with chroma/hue preserved.
+The default scaling is computed in GenSpace because the gradient-uniformity metrics dominate authoring quality.
 Dark/light mode.
-When surround parameters are trained, the surround $S$ directly modulates the transform.
-Currently, a fallback soft L-inversion maps colors between light ($S=0.7$) and dark ($S=0.2$) modes with contrast enforcement.
+The surround parameter $S$ is architecturally present but in v21 is held at zero.
+The current dark/light mode adaptation uses a soft L inversion across the GenSpace L axis, with contrast enforcement on the resulting pair.
+A future MetricSpace release with trained $S$ parameters will replace this heuristic.
 Token export.
-A single TokenExporter produces CSS custom properties (oklch(), color(display-p3 ...)), Android XML (using HCT tone/chroma/hue semantics for Material Design compatibility), iOS Swift (Display P3), Tailwind config, and raw JSON.
+A single TokenExporter produces CSS custom properties (oklch(), color(display-p3 ...), color(--helmgenlch ...)), Android XML (using HCT tone/chroma/hue semantics for Material Design compatibility), iOS Swift (Display P3), Tailwind config, and raw JSON.
+Each target reads from the same in-memory token tree, so a token edit propagates to every output format simultaneously.
 9 Limitations and Future Work
@@
     Model complexity.
-72 parameters is considerably more than Oklab ($\sim$15 designed parameters).
+72 trained parameters is considerably more than Oklab ($\sim$15 designed parameters).
 Each stage has a perceptual motivation, but the overall model is harder to interpret than simpler spaces.
+The architecture additionally reserves 22 surround/hue-modulated slots that are held at zero in v21.
   3. 3.
-    Hue alignment. RMS hue error of 18.1° is a compromise. Some hue regions (cyan) have errors up to 23.9°. Further harmonic terms could improve this at the risk of overfitting.
+[…]
   4. 4.
-    Surround dependency. The surround parameter $S$ is architecturally present but not yet trained on viewing-condition data. Current dark/light adaptation uses heuristics.
+    Hue alignment.
+RMS hue error of 26.4° is a substantial compromise; the worst primary (cyan) is 53.4° off the HSL reference.
+v21’s training pushed parameters into the distance-optimal basin at the cost of geometric uniformity — a trade we accept because the distance metric is the headline product, but applications that need accurate hue specification should compose MetricSpace with an explicit hue-rotation post-step or use GenSpace (where the geometry is preserved by construction).
   5. 5.
+    Surround dependency.
+The surround parameter $S$ is architecturally present but in v21 the 11 surround-suffix parameters are all held at zero (training data is average-surround only).
+Current dark/light adaptation uses a heuristic L-inversion in GenSpace with contrast enforcement.
+  6. 6.
     Sub-dataset variation.
-Helmlab underperforms CIEDE2000 on 3 of 6 combvd sub-datasets (Table˜4).
-The blue-band refit improves 5 of 6 sub-datasets compared to the previous version; only BFD-P(C), measured under illuminant C with a 2° observer, remains notably worse (30.15 vs 30.05).
-  6. 6.
-    Not general-purpose. Helmlab is optimized for UI workflows (sRGB/P3 gamut, screen viewing). It is not designed for print, photography, or spectral applications.
+v21 trails CIEDE2000 on three of six combvd sub-datasets (LEEDS, RIT-DuPont, WITT).
+The lead on BFD-P(D65) and BFD-P(M) is large enough to dominate the aggregate, but designers comparing spaces on a single small sub-dataset may pick different winners.
   7. 7.
-    Blue-band trade-off.
-The blue-band refit (Section˜6.6) reduces gradient non-uniformity by $8.9\times$ at a cost of +0.08+0.08 stress on combvd.
-The Munsell CV increases from 17.6% to 20.0%, hue RMS from 16.1° to 18.1°, and the He et al.
-cross-validation stress increases from 29.0 to 30.3.
+    Not general-purpose.
+Helmlab is optimized for UI workflows (sRGB/P3 gamut, screen viewing, D65 illuminant).
+It is not designed for print, photography, illuminant-A scenes, or spectral applications.
+CAM16 remains the appropriate choice for arbitrary viewing conditions.
   8. 8.
@@
 Bootstrap confidence intervals confirm the aggregate improvement over CIEDE2000 ($p<10^{-4}$), but per-sub-dataset significance and parameter uncertainty across restarts remain unquantified.
-The open-source release includes a second, generation-optimized space (GenSpace) with shared $\gamma\!=\!\tfrac{1}{3}$ and no enrichment stages, which trades distance accuracy for faster, invertible gradient and palette generation.
-For future work, we plan to train the surround parameter on paired light/dark viewing data and to expand the training set with targeted human feedback using the bidirectional optimization framework that is already in the codebase.
+  9. 9.
+    GenSpace sRGB round-trip.
+GenSpace’s sRGB round-trip is $\sim 5.6\!\times\!10^{-8}$, six orders of magnitude looser than MetricSpace’s.
+This is below the visual threshold and matches OKLab’s worst-case round-trip in the same regions, but it is looser than the other GenSpace round-trips (P3, Rec.
+2020 are at machine epsilon).
+The cause is the L-gated enrichment Newton inversion in the blue band; an analytic inverse for the enrichment is open work.
+For future work, we plan to train the surround parameters on paired light/dark viewing data, to add an analytic enrichment inverse for GenSpace, and to expand the training set with targeted human feedback using the bidirectional optimisation framework already present in the codebase.
 10 Conclusion
-We have presented Helmlab, a 72-parameter analytical color space that achieves stress 23.30 on combvd, which is 20.2% lower than CIEDE2000.
-The achromatic axis is guaranteed to have $C<10^{-6}$ via the neutral correction, and a rigid rotation brings hue error down to 18.1° RMS without affecting the distance metric.
-A blue-band refit with sub-dataset penalties reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ at a cost of only +0.08+0.08 stress points, while improving stress on 5 of 6 sub-datasets.
-The neutral correction is the key element that makes the space usable for both measurement and generation: it lets the optimizer find its best parameters freely and then fixes the achromatic axis.
-Code, trained parameters, and utilities are available at https://github.com/Grkmyldz148/helmlab.
+[…]
+The neutral correction is the architectural element that lets the two specialisations co-exist: MetricSpace can chase the distance-fit basin freely, because the achromatic axis is then restored exactly by post-correction; GenSpace achieves the same guarantee structurally, by construction.
+Together with the rigid rotation invariance, this lets us decouple “what is the best distance metric” from “what should the chroma plane look like for generation”.
+Helmlab is shipped in production: helmlab on PyPI and npm, four spaces and a $\Delta E$ method merged into Color.js (PR #722), and a PostCSS plugin (postcss-helmlab) for build-time CSS function transformation.
+Source, trained parameters, and a full ColorBench-driven benchmark report are at https://github.com/Grkmyldz148/helmlab.
 An interactive demo is at https://grkmyldz148.github.io/helmlab/demo.html and documentation at https://grkmyldz148.github.io/helmlab/.
 Acknowledgments
-Thanks to the 71 anonymous observers who participated in the bidirectional evaluation study during development.
-Appendix A Parameter Table
-Table˜5 lists all 72 optimized parameters grouped by pipeline stage.
-Table 5: Optimized Helmlab parameters (v20b, $\varphi=-28.2$°).
+We thank the 71 observers — listed by first name or chosen handle in the released dataset (datasets/human_feedback.json, github.com/Grkmyldz148/helmlab) — who contributed 3,552 paired-comparison judgments during development.
+We also gratefully acknowledge the open-source review of the underlying PR #722 on color-js/color.js: in particular @facelessuser (near-black achromatic instability, M2 normalisation regressions, the clean 64-bit matrix derivation methodology), @lloydk (the WebGPU chart library that surfaced the h $\approx$ 264° blue fold and related cubic-cusp artefacts), @Userminusone (the Metal-shader hue/ lightness probe, scale-invariance critique of NR, the boundary-warp straightening alternative), and @svgeesus (the original paper proposal, the “0% should be achromatic” authoring convention, and the ICC Displays working-group framing).
+The two-month review thread was, in effect, a parallel research process; this paper is its written record.
+Appendix A Parameter Table (MetricSpace v21)
+Table˜6 lists all 72 trained MetricSpace v21 parameters grouped by pipeline stage.
+The architecturally-reserved 22 surround/hue-modulated parameters are held at zero in v21 and omitted from this table; their full list is in the JSON checkpoint.
+Table 6: MetricSpace v21 parameters ($\varphi=-28.2$°). Trained on combvd + He 2022 + MacAdam 1974 with sub-dataset-balanced loss (Equation˜16).
 Stage Parameter Value Description
-$\mathbf{M}_{1}$ (9) Row 0 [0.734, 0.240, -0.158] XYZ $\to$ cone-like
-Row 1 [-0.330, 1.235, -0.000]
-Row 2 [0.086, 0.341, 0.832]
-$\gamma$ (3) $\gamma_{0},\gamma_{1},\gamma_{2}$ 0.389, 0.416, 0.424 Power compression
-$\mathbf{M}_{2}$ (9) Row 0 [-0.420, 0.485, 0.642] Compressed $\to$ Lab
-Row 1 [1.933, -2.718, 0.764]
-Row 2 [0.006, 1.628, -1.261]
-Hue corr. (8) $\alpha_{1..4},\beta_{1..4}$ see JSON 4-harmonic Fourier
-H-K (6) $w,p,m,s_{1},c_{2},s_{2}$ 0.389, 0.849, -0.494, … Chroma $\to$ L
-Light. (8) $p_{1},p_{2},p_{3},Lh_{c1},Lh_{s1},\lambda_{d},h_{c},h_{s}$ 0.222, -0.727, 0.695, … Cubic + Lh + dark
-Chroma (18) 4-harm + power + $lc_{1},lc_{2}$ + HLC see JSON Hue/power/L-dep
-Hue-L (4) $g_{c1..2},g_{s1..2}$ see JSON Hue $\to$ L
-Distance (7) $s_{L}$ $1.01\times 10^{-3}$ Pair-dep L weight
-$s_{C}$ 0.022 Pair-dep C weight
-pp 0.804 Minkowski exponent
-$w_{C}$ 1.046 Chroma weight
-$c$ 1.590 Compression
-$q$ 1.100 Post-power
-$\alpha$ 0.000 Linear term
-Full parameter values are available in the accompanying metric_params.json file.
+$\mathbf{M}_{1}$ (9) Row 0 [-0.7213, -0.4534, -0.1929] XYZ $\to$ cone-like
+Row 1 [-0.7882, -1.7952, -0.0876]
+Row 2 [-0.0918, -0.4577, -1.2922]
+$\gamma$ (3) $\gamma_{0},\gamma_{1},\gamma_{2}$ 0.4723, 0.5149, 0.5113 Signed power compression
+$\mathbf{M}_{2}$ (9) Row 0 [-0.2636, -0.4168, -0.4927] Compressed $\to$ Lab
+Row 1 [-1.8898, -3.1212, -1.0422]
+Row 2 [-0.3585, -1.7694, -1.4121]
+Hue corr. (8) cos1..4, sin1..4 see JSON 4-harmonic Fourier rotation $\delta(h)$
+H-K (6) $w_{\mathrm{HK}},p_{\mathrm{HK}},\mathrm{hue\_mod}$ 0.2676, 0.8935, 0.7173 Base + envelope
+sin1, cos2, sin2 0.6915, 0.4865, 0.9853 2-harmonic hue mod $f_{\mathrm{HK}}(h)$
+Lightness (8) $p_{1},p_{2},p_{3}$ 0.5385, 0.1251, 0.6769 Cubic correction
+$Lh_{c1},Lh_{s1}$ -0.4963, -0.0956 Hue-modulated additive
+$\lambda_{d},h_{c},h_{s}$ -0.0291, 1.3347, -0.1699 Dark-region exp compression
+Chroma (18) cs1..4, cp1..2, $lc_{1,2}$, hlc1,2 see JSON 4-harm CS + 2-harm CP + L-dep + HLC
+Hue-L (4) $g_{c1..2},g_{s1..2}$ see JSON 2-harmonic $L\mathrel{*}\!\!=\!\exp(g(h))$
+Distance (7) $s_{L}$ -0.9155 Pair-dep L weight (Equation˜12)
+$s_{C}$ -2.9268 Pair-dep C weight (Equation˜13)
+$p$ -1.9737 Minkowski exponent
+$w_{C}$ -3.9660 Chroma weight
+$c$ -52.473 Compression
+$q$ -0.4790 Post-compress power
+$\alpha$ -0.0000 Linear-asymptote term (unused in v21)
+Full parameter values, including the architecturally-reserved zero slots and all chroma-stage Fourier coefficients, are available in research/checkpoints/metricspace_v21.json.
+Appendix B Parameter Table (GenSpace v0.11.1)
+Table 7: GenSpace v0.11.1 parameters.
+Stage Value
+$\mathbf{M}_{1}$ $\times3$, see JSON (Bradford CAT pre-baked)
+Transfer depcubic, $\alpha=0.021$
+$\mathbf{M}_{2}$ $\times3$, see JSON
+Chroma power 0.978
+PW $L$ correction 19-point dense, see JSON
+Enrichment L-gated hue rotation
+amplitude 0.058
+center hue 264.5°
+$\sigma$ 0.7 (rad)
+gate $L$ window [0.37, 1.0]
+Full parameter values ship bundled with the helmlab package (helmlab.GenSpace().params) and are written out as src/helmlab/data/gen_params.json in the source distribution.
+White(D65) maps to $L=1$ exactly, the achromatic axis is held at $|a|,|b|<10^{-4}$, and the structurally-achromatic property holds because the depcubic transfer is signed-symmetric and the smooth neutral-blend in Stage 2 collapses sub-microscopic channel spread to the channel mean before $\mathbf{M}_{2}$ projection.